### v2

# LAGUNA: LAnguage Guided UNsupervised Adaptation with structured spaces

Abstract
Unsupervised domain adaptation remains a critical challenge in enabling the knowledge transfer of models across unseen domains. Existing methods struggle to balance the need for domain-invariant representations with preserving domain-specific features, which is often due to alignment approaches that impose the projection of samples with similar semantics close in the latent space despite their drastic domain differences. We introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach that shifts the focus from aligning representations in absolute coordinates to aligning the relative positioning of equivalent concepts in latent spaces. LAGUNA defines a domain-agnostic structure upon the semantic/geometric relationships between class labels in language space and guides adaptation, ensuring that the organization of samples in visual space reflects reference inter-class relationships while preserving domain-specific characteristics. Remarkably, LAGUNA surpasses previous works in $18$ different adaptation scenarios across four diverse image and video datasets with average accuracy improvements of $+3.32\%$ on DomainNet, $+5.75\%$ in GeoPlaces, $+4.77\%$ on GeoImnet, and $+1.94\%$ mean class accuracy improvement on EgoExo4D.
1 Introduction
Domain shift challenges trained models to generalize across scenarios with differing distributions and presents a significant hurdle for supervised learning in computer vision [42]. While fine-tuning with labeled data from the target domain can mitigate this problem, obtaining such labels often proves prohibitive. Unsupervised domain adaptation (UDA) [4] offers a compelling alternative, enabling knowledge transfer to novel domains without relying on expensive labeled data, which has garnered significant attention [70], promising cost-effective solutions for real-world applications prone to domain shift.
Figure 1: Top: Existing Unsupervised Domain Adaptation approaches align source and target spaces in absolute coordinates, which may overlook the peculiarities of the distinct domains resulting in partial alignment. Bottom: LAGUNA aligns source (left) and target (right) spaces in relative terms, allowing data points to be distinct in absolute coordinates (e.g., circles in source and target), while $\theta_{i}^{s}$, $\theta_{i}^{t}$, their angles with other data points in the original spaces, match $\theta_{i}^{l}$, those of the reference domain-agnostic structure derived from language ($\theta_{i}^{s}$ $\sim\theta_{i}^{l}\sim$ $\theta_{i}^{t}$), encouraging the development of similar geometric-semantic relations.
The typical UDA setting considers the availability of a labeled source domain and an unlabeled target domain. In general, source and target domains are semantically equivalent but drawn from distinct data distributions, e.g., real images versus clipart of chairs, TVs, and mugs. Thus, the main challenge of UDA is to effectively mitigate the distribution shift between domains, which is often addressed by reducing the distribution discrepancy of source and target representation spaces either by minimizing some discrepancy measure [21, 58, 63], using adversarial learning [68, 67, 33, 77, 49, 31, 40], aligning data around centroids [33], or leveraging pseudo-labels [27, 33]. These methods aim to align source and target representation spaces in a shared coordinate system, pushing feature vectors of equivalent semantic concepts close to each other in the embedding space, which may happen at the expense of the representation power of individual domain-specific spaces. For instance, bright colors and rounded shapes might be important to encode for representing a clipart, while nuanced shadows, reflections, and textures might be important to represent a real image. As a result, the aligned space may become overly generic, correctly encoding only a subset of the data (see Figure 1(top)). Recent work [46] showed that semantically equivariant representation spaces of similarly trained neural networks exhibit distinct representation spaces with matching geometrical structures. For instance, two data points $(x_{1},x_{2})$ may be mapped to distinct vector pairs in the two representation spaces $(v_{1}^{1},v_{2}^{1})$ and $(v_{1}^{2},v_{2}^{2})$ (e.g., $||v_{1}^{1}-v_{1}^{2}||_{2}>>0$ and $||v_{2}^{1}-v_{2}^{2}||_{2}>>0$), while angles between the two pairs will be similar in their own representation spaces (e.g., $\angle(v_{1}^{1},v_{2}^{1})\sim\angle(v_{1}^{2},v_{2}^{2})$) [^1] The reader is referred to [46] for empirical verifications of such claim.. This suggests that pushing representation spaces to overlap in absolute coordinates, as done in current domain adaptation approaches, is not necessary to obtain equivariant representation spaces.
Based on this observation, we introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach to domain adaptation which guides source and target spaces to develop semantic-geometric inter-relationships reflecting the structure of a reference space (Figure 1(bottom)). As shown in recent works, language can provide a semantic space agnostic to the nuances of visual observations enabling robust zero-shot generalization [54, 10] and supporting domain robustness [14, 66, 23, 45, 27], hence we choose language to build our reference structured space in LAGUNA. This assumes the availability for both source and target samples of natural language descriptions, which can be generated from available captioning models [32] or collected from the web at a fraction of the cost of typical human labeling procedures [27]. Specifically, LAGUNA employs a 3-stage approach to structurally align the source and target representation spaces while allowing each domain to preserve its typical patterns. In Stage 1, textual class labels are mapped to a domain-agnostic reference space representing their relationships at the conceptual level. In Stage 2, a language model is trained to map textual captions to the reference latent space in order to provide pseudo-labels for target samples Finally, in Stage 3, a cross-domain classifier is trained to encourage domain-specific representations to follow the reference structure.
Experiments demonstrate the superiority of LAGUNA over existing state-of-the-art methods across $18$ different domain adaptation scenarios sourced from four diverse image and video datasets, with gains of $+3.32\%$ on DomanNet [50], $+5.75\%$ on GeoPlaces [26], $+4.77\%$ on GeoImnet [26] and $+1.94\%$ mean per class accuracy on EgoExo4D [20]. We further report ablations to analyze the specific contributions of our design choices.
In sum, our main contributions are as follows: 1) we investigate the use of relative representations for UDA, showing its advantages with respect to the current absolute alignment paradigm; 2) we propose LAGUNA, a three-stage method which learns a cross-domain classifier where individual source and target spaces are distinct yet aligned to a reference structure derived from language; 3) through extensive ablations and comparisons with state-of-the-art models, we show the superiority of the proposed approach. Code available at: https://github.com/ADiko1997/LAGUNA.git
2 Related Work
2.1 UDA in Computer Vision
UDA seeks to transfer knowledge from a labeled source domain to an unlabeled target domain [2, 16, 40]. This is tackled through various approaches, prominently, discrepancy-based methods that minimize distribution differences using techniques like MMD [63, 39, 61, 28, 58, 31], and adversarial learning [4, 64, 57, 6, 68, 67, 40, 16, 64, 38]. Other approaches also investigated class-conditional distributions alignment [49, 43, 72, 29, 68], clustering similar instances across domains [9, 48, 36, 24], instance-specific adaptation[60, 25, 65], and self-training leveraging pseudo-labeling to improve target domain performance [15, 36, 62, 52, 27]. More recently, transformer-based approaches utilize patch composition for domain-invariant representations through intermediate domains [77]. Video domain adaptation tackles the unique challenges of temporal dynamics and consistency in videos [8, 69, 56] with a particular focus on adapting models across the exocentric and egocentric domains [53, 22, 34, 74]. Existing approaches sought to align source and target representation spaces in absolute coordinates, often falling short in bridging complicated forms of domain shift [26, 52]. Contrarily, LAGUNA aligns representations in relative terms, encouraging source and target visual spaces to share a similar structure as a reference space derived from language while allowing them to encode domain-specific peculiarities.
Figure 2: The 3 stages of LAGUNA. In Stage 1, a domain-agnostic semantic structure in the form of a set of anchors $\mathcal{A}$ is defined. In Stage 2, a language model $\mathcal{G}$ is trained on source captions-label pairs to provide pseudo-labels for the target domain. The model is trained to be structurally aligned to anchors ${}^{*}\mathcal{A}$ with a structural loss $\mathcal{L}_{S}$ and to predict class labels with a cross-entropy loss $\mathcal{L}_{CE}$. In Stage 3, a cross-domain visual classifier is trained on labeled target and unlabeled source images. A visual encoder $\mathcal{V}$ is used to represent target and source samples as $g_{i}^{t}$ and $g_{i}^{s}$. The cosine similarities $\mathbf{r}^{g_{i}^{t}}$ between the target visual representation $g_{i}^{t}$ and all learnable target anchors $\mathcal{A}_{t}$ are bound to be similar to the cosine similarities $\mathbf{r}^{z_{i}^{t}}$ between the textual representation of the target caption $z_{i}^{t}$ predicted through the language model $\mathcal{G}$ and the reference domain-agnostic anchors ${}^{*}\mathcal{A}$ by means of the structural loss $\mathcal{L}_{S}$. A similar processing is applied to the visual representation of source sample $g_{i}^{s}$, the textual representation of the ground truth class ${}^{*}A[y_{i}^{S}]$ and the learnable source anchors $\mathcal{A}_{s}$. A regularization loss $\mathcal{L}_{Reg}$ is used to avoid collapse during anchor learning. Representations $g_{i}^{t}$ and $g_{i}^{s}$ are hence grounded into the learnable source anchors $\mathcal{A}_{s}$ through a novel Cross-Domain Attention layer taking visual representations as queries, source/target anchors as keys and source anchors as values. Classification is finally achieved through an MLP classifier supervised through a cross-entropy loss $\mathcal{L}_{CE}$ using ground truth labels in the source domain and pseudo-labels in the target domain. The visual encoder, Cross-Domain Attention, and MLP layers share weights across domains.
2.2 Language Guidance in Vision UDA
Vision-language models like CLIP [54] have shown promising zero-shot transfer capabilities [10]. However, when fine-tuned to a specific scenario [51, 1], they often lose the ability to generalize to new domains [30, 71]. To address this issue, recent works leveraged textual information to bridge the gap between domains [14, 66, 23, 45, 18, 19]. In particular, the seminal work of [27] proposed to use rich textual captions to provide pseudo-labels for the target domain and addressed UDA by training a joint classifier on source and target domains. Similarly to [27], we generate pseudo-labels from textual captions, but, rather than seeking to align the visual representation spaces of source and target domains to language in an absolute reference frame, LAGUNA uses language to guide a relative alignment between source and target visual spaces.
2.3 Relative Encodings
Recent work [46] has shown that equivalent latent spaces of similarly trained networks tend to be misaligned in absolute terms but share similar internal geometric relationships. As a result, representing data points with relative encodings, obtained as vectors of cosine similarity values of a data point with a predefined set of anchors, allows to perform zero-shot model stitching [46], translate representation spaces across models [44] or modalities [47]. As shown in [5], predefined invariances can be incorporated into the learned representation to enable specific forms of relative representations. In [11] relative encodings were used to tackle the downstream task of action anticipation. LAGUNA builds on relative encodings to represent the semantic inter-relations between classes in a domain-agnostic language-based reference space and support the development of aligned yet specialized source and target domains.
3 Method
In Stage 1, a language-based reference structure space is defined. We represent this structure as a set of anchors, which we will use as a reference to guide representation learning. In Stage 2, a text-supervisor language model is trained to map image/video captions to class categories in order to provide pseudo-labels for the unlabeled target domain. The language model is also trained to keep the internal text embeddings aligned to the domain-agnostic structure defined in Stage 1. Finally, in Stage 3, a cross-domain visual classifier is trained to categorize source and target samples by learning domain-specific action anchors bound to be structurally similar to the domain-agnostic structure defined in Stage 1.
3.1 Problem Setup
We follow the formulation of [27] and assume a labeled source domain $\mathcal{X}_{s}:\{(x_{i}^{s},y_{i}^{s},l_{i}^{s})\}_{i=1}^{N_{s}}$, where samples $x_{i}^{s}$ are paired both with labels $y_{i}^{s}$ and language descriptions $l_{i}^{s}$, whereas the target domain $\mathcal{X}_{t}:\{x_{i}^{t},l_{i}^{t}\}_{i=1}^{N_{t}}$ contains unlabeled samples $x_{i}^{t}$ paired with language descriptions $l_{i}^{t}$. These textual descriptions can be readily obtained from associated metadata or generated using image-to-text models [32]. $N_{s}$ and $N_{t}$ represent the number of source and target samples, respectively, and the two domains share the same high-level semantics and number of classes $N_{c}$.
3.2 Stage 1 - Reference Structure Definition
Assuming shared classes across domains, in Stage 1 (Fig. 2), we construct a language-based, domain-agnostic reference structure in the form of a set of vectors $\mathcal{A}\in\mathbb{R}^{N_{c}\times D_{l}}$. Each vector is obtained by encoding the $N_{c}$ class label names through a pre-trained SentenceTransformer model with output dimensionality $D_{l}$ trained for semantic similarity [55]. Following literature on relative representations [46], we treat the vectors in $\mathcal{A}$ as a reference set of anchors. Given a vector $v\in\mathbb{R}^{D_{l}}$, we define its relative encoding with respect to anchors $\mathcal{A}$ as $\mathbf{r}^{v}=rel(v,\mathcal{A})=[\cos(v,\mathcal{A}[1]),\ldots,\cos(v,\mathcal{A}[N_{c}])]$, where $\cos(\cdot,\cdot)$ is the cosine similarity (i.e., the cosine of the angle $\theta$ between two vectors), and $\mathcal{A}[i]$ is the anchor in $\mathcal{A}$ associated to class $i$. Intuitively, vector $\mathbf{r}^{v}$ encodes the geometrical/semantic relationships between vector $v$ and the reference language anchors $\mathcal{A}$. Following this logic, the set of all reference affinities is defined as $\mathbf{r}^{\mathcal{A}}=[rel(\mathcal{A}[i],\mathcal{A}),\dots,rel(\mathcal{A}[N_{c}],\mathcal{A})]$ where each class anchors $\mathcal{A}[i]$ is represented according to its relationship with respect to the other anchors in $\mathcal{A}$. During the training of the cross-domain visual classifier (Stage 3), these encodings are used to enforce the learned latent space to follow a similar structure as the one induced by $\mathcal{A}$.
3.3 Stage 2: Training of the Language Supervisor
Similar to [27], we train a language model to map image captions to semantic labels, with the aim to provide pseudo-labels for target samples, which are associated with captions but not class labels. Differently from [27], we use the language supervisor both to provide pseudo-labels for target samples, useful to train the classifier, and textual representations semantically structured as the domain-agnostic anchors $\mathcal{A}$, useful to encourage alignment to the reference structure. Hence, rather than training a regular classifier, we directly supervise a language model $\mathcal{G}$ to provide representations that are 1) geometrically aligned to $\mathcal{A}$ and 2) suitable for predicting class labels. Specifically, given a pair of source caption and corresponding label $(l_{i}^{s},y_{i}^{s})$, $\mathcal{G}$ processes $l_{i}^{s}$ to produce a vector representation $z^{s}_{i}$:
$z_{i}^{s}=\mathcal{G}(l_{i}^{s}).$ (1)
Next, we encourage the vector representation $z_{i}^{s}$ to be geometrically aligned to the anchor $\mathcal{A}[y_{i}^{s}]$ corresponding to ground truth action $y_{i}^{s}$. To do so, we compute $\mathbf{r}^{z_{i}^{s}}=rel(z_{i}^{s},\mathcal{A})$, the relative encoding of $z_{i}^{s}$, and supervise it to be similar to $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$, the relative encoding of the anchor $\mathcal{A}[y_{i}^{s}]$, with the following structure-preserving loss:
$\mathcal{L}_{S}=|\mathbf{r}^{z_{i}^{s}}-\mathbf{r}^{y_{i}^{s}}|.$ (2)
This loss encourages the encodings $z_{i}^{s}$ of text descriptions $l_{i}^{s}$ associated with labels $y_{i}^{s}$ to preserve the same geometrical associations as their corresponding anchor, encouraging the latent space learned by $\mathcal{G}$ to mirror the structure defined by $\mathcal{A}$. To further favor alignment to $\mathcal{A}$, rather than employing a classification head, we predict class probabilities directly by Softmax-normalizing relative encodings:
$p^{z_{i}^{s}}_{j}=\frac{e^{\mathbf{r}^{z_{i}^{s}}_{j}}}{\sum_{k}e^{\mathbf{r}^{z_{i}^{s}}_{k}}}.$ (3)
Finally, we train the model using a combined loss $\mathcal{L}$:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(p^{z_{i}^{s}},y_{i}^{s})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{z_{i}^{s}},\mathbf{r}^{y_{i}^{s}}),$ (4)
where $\mathcal{L}_{CE}$ is the cross-entropy loss, while $\lambda_{1}$ and $\lambda_{2}$ are hyperparameter weights to calibrate the magnitude of each loss. We refer to the label predicted by $\mathcal{G}$ from $l_{i}^{t}$ as $\overline{y}_{i}^{t}$.
3.4 Stage 3: Cross-Domain Visual Classifier
Stage 3 aims to train a cross-domain visual classifier aligning representations extracted through a visual encoder $\mathcal{V}$ to the structure imposed by the domain-agnostic language anchors $\mathcal{A}$. We allow each domain to develop its own latent space, but we encourage both spaces to be aligned to two sets of learnable anchors $\mathcal{A}_{t}\in\mathbb{R}^{N_{c}\times D_{v}}$ and $\mathcal{A}_{s}\in\mathbb{R}^{N_{c}\times D_{v}}$ specific to the target and source domain respectively. We further ensure that the structures of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are aligned to the domain-agnostic, fixed anchors $\mathcal{A}$ through the supervision provided by ground truth labels $y_{i}^{s}$ in the source domain and the text supervision provided by $\mathcal{G}$ in the target domain in the form of the predicted vector $z_{i}^{t}$ and the predicted pseudo-label $\overline{y}_{i}^{t}$. The source and target datasets are merged ($\hat{\mathcal{X}}=\mathcal{X}_{t}+\mathcal{X}_{s}$) and used for training with a total of $M=N_{s}+N_{t}$ samples. Source samples comprise an image, caption, and label triplet $(x_{i}^{s},l_{i}^{s},y_{i}^{s})$, whereas target samples $(x_{i}^{t},l_{i}^{t},\overline{y}_{i}^{t})$ replace label $y_{i}^{s}$ with the pseudo-label $\overline{y}_{i}^{t}$. Target and source images are processed by $\mathcal{V}$ to obtain latent representations $g_{i}^{t}$ and $g_{i}^{s}$, while target captions $l_{i}^{t}$ are processed by $\mathcal{G}$ to obtain latent representations $z_{i}^{t}$ as follows:
$g_{i}^{t}=\mathcal{V}(x_{i}^{t}),\hfill z_{i}^{t}=\mathcal{G}(l_{i}^{t}),\hfill g_{i}^{s}=\mathcal{V}(x_{i}^{s}).$ (5)
These representations and the anchors are used to achieve two main training objectives: 1) structure learning and 2) classification training.
3.4.1 Structure Learning
To ensure the learned visual representations adhere to the structure defined by $\mathcal{A}$, LAGUNA employs a supervised learning approach based on relative encodings. For target domain samples (without ground truth labels), we first compute the relative encoding $\mathbf{r}^{z_{i}^{t}}=rel(z_{i}^{t},\mathcal{A})$. For source domain samples, we compute the relative encoding of the anchor associated to the ground truth class $\mathcal{A}[y_{i}^{s}]$: $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$. These encodings represent the relative positions of textual representations with respect to the reference space $\mathcal{A}$ and are not trainable as $\mathcal{G}$ is fixed. On the visual side, we compute the relative encodings of $g_{i}^{t}$ and $g_{i}^{s}$ with respect to the learnable anchors $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$, i.e., $\mathbf{r}^{g_{i}^{t}}=rel(g_{i}^{t},\mathcal{A}_{t})$ and $\mathbf{r}^{g_{i}^{s}}=rel(g_{i}^{s},\mathcal{A}_{s})$. We encourage the visual relative encodings ($\mathbf{r}^{g_{i}^{t}},\mathbf{r}^{g_{i}^{s}}$) to match the text relative encodings ($\mathbf{r}^{z_{i}^{t}},\mathbf{r}^{y_{i}^{s}}$) with an L1 loss:
$\mathcal{L}_{S}=\begin{cases}|\mathbf{r}^{g_{i}^{t}}-\mathbf{r}^{z_{i}^{t}}|,&\textit{if Target Domain}\\ |\mathbf{r}^{g_{i}^{s}}-\mathbf{r}^{y_{i}^{s}}|,&\textit{otherwise}\end{cases}.$ (6)
This loss encourages the visual representations to maintain relationships with their respective domain-specific visual anchors that mirror the relationships expressed by the domain-invariant anchors $\mathcal{A}$. This process supervises both the learnable anchors $\mathcal{A}_{s},\mathcal{A}_{t}$ and the visual encoder $\mathcal{V}$. However, since this loss alone might lead to the collapse of anchor representations into a small region, hindering effective classification, we introduce a volume (or spread) regularization loss. This loss treats each set of anchors as a multidimensional parallelotope whose spread or volume occupied in the latent space can be measured by the determinant of its Gram matrix. Given the Gram matrices for the three sets of anchors:
$\gamma=\mathcal{A}\mathcal{A}^{T},\hfill\gamma_{s}=\mathcal{A}_{s}\mathcal{A}_{s}^{T},\hfill\gamma_{t}=\mathcal{A}_{t}\mathcal{A}_{t}^{T},$ (7)
we encourage the volume of each domain-specific parallelotope to be approximately equal to the volume of the reference anchors $\mathcal{A}$ with the following regularization loss: e
$\mathcal{L}_{Reg}$ $=$ $|logDet(\gamma_{t})-logDet(\gamma)|+$ (8)
$|logDet(\gamma_{s})-logDet(\gamma)|,$
where we consider log-determinants instead of determinants for numerical stability. This regularization loss ensures the visual anchors occupy a volume similar to the one of $\mathcal{A}$, preventing collapse and promoting better separation of classes in the representation space.
3.4.2 Classification Training
The network undergoes classification training concurrently with structure learning. To ensure that source and target visual representations are grounded in the structure imposed by $\mathcal{A}$, while still being free to capture domain-specific nuances, we propose a novel cross-domain attention layer that aims to ground domain-specific visual representations $g_{i}^{t}$ and $g_{i}^{s}$ to the structure of source anchors $\mathcal{A}_{s}$. We include two Cross-Domain attention layers sharing the same weights for the target and source branches of the architecture (See Fig. 2). In the target branch, we use $g_{i}^{t}$ as queries, target anchors $\mathcal{A}_{t}$ as keys, and source anchors $\mathcal{A}_{s}$ as values. The output is summed to $g_{i}^{t}$ with a residual connection to include both domain-specific and cross-domain information in the final representation:
$f_{i}^{t}=\text{Attention}(Q=g_{i}^{t},K=\mathcal{A}_{t},V=\mathcal{A}_{s})+g_{i}^{t}$ (9)
A similar processing is applied to the source encoding $g_{i}^{s}$:
$f_{i}^{s}=\text{Attention}(Q=g_{i}^{s},K=\mathcal{A}_{s},V=\mathcal{A}_{s})+g_{i}^{s}$ (10)
Crucially, the cross-domain attention layer constructs an attention map by leveraging the domain-specific relations between the encoded representations and the anchors (i.e., $QK^{T}$). These relations are expected to be similar across domains due to the influence of the structural loss $\mathcal{L}_{S}$. This process results in a unified representation since the output of the attention layer always depends on values coming from the source anchors $\mathcal{A}_{s}$. The residual connection included in the attention layer is introduced to account for domain-dependent class characteristics, which results in final feature vectors $f_{i}^{s}$ and $f_{i}^{t}$. Finally, an MLP classification head processes $f_{i}^{t}$ and $f_{i}^{s}$ to output predictions $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{s}$. The MLP has shared weights across domains and is optimized with a cross-entropy loss using the ground truth label $y_{i}^{s}$ for source examples and the pseudo-label $\overline{y}_{i}^{t}$ predicted by $\mathcal{G}$ in the case of target samples. The overall training objective of the classifier is the following:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(\hat{y_{i}},y_{i}^{*})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{*},\mathbf{r}^{g_{i}})+\lambda_{3}\mathcal{L}_{Reg}(\gamma,\gamma_{*}),$ (11)
where $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ are hyperparameter weights to calibrate the magnitude of each loss value, and ’*’ signifies that the argument is domain-dependent [^2] For example, for $\mathcal{L}_{CE}$ the ground truth can either be $y_{i}^{s}$ or $\overline{y}_{i}^{t}$..
4 Experiments
[…]
Source Only 63.02 49.47 60.48 70.52 56.09 52.53 70.42 65.91 54.47 73.34 60.09 48.25 60.38
MCD [58] 39.40 25.20 41.20 44.60 31.20 25.50 34.50 37.30 27.20 48.10 31.10 22.80 34.01
MDD [75] 52.80 41.20 47.80 52.50 42.10 40.70 54.20 54.30 43.10 51.20 43.70 41.70 47.11
CGDM [13] 49.40 38.20 47.20 53.50 36.90 35.30 55.60 50.10 43.70 59.40 37.70 33.50 45.04
SCDA [33] 54.00 42.50 51.90 55.00 44.10 39.30 53.20 55.60 44.70 56.20 44.10 42.00 48.55
SSRT-B [62] 69.90 58.90 66.00 75.80 59.80 60.80 73.20 70.60 62.20 71.40 61.70 55.20 65.41
MemSAC [25] 63.49 42.14 60.32 72.33 54.92 46.14 73.46 68.04 52.75 74.42 57.79 43.57 59.11
CDTrans [73] 66.20 52.90 61.50 72.60 58.10 57.10 72.50 69.00 59.00 72.10 62.90 53.90 63.16
PMTrans [77] 74.10 61.10 70.00 79.30 63.70 62.70 77.50 73.80 62.60 79.80 69.70 61.20 69.63
UDA with Language Guidance
TextMatch [27] 71.36 64.30 65.32 81.25 65.65 64.85 81.09 72.65 63.94 81.08 70.84 64.17 70.64
nGramMatch [27] 68.92 59.82 63.15 76.35 61.72 62.87 76.35 69.28 62.51 76.04 68.52 60.52 67.17
LaGTran [27] 77.30 68.25 67.35 81.31 67.03 66.81 80.78 75.62 68.08 79.23 73.80 63.44 72.41
LAGUNA 80.34 70.68 71.92 83.07 69.51 70.59 83.34 79.71 70.51 83.32 77.47 68.32 75.73
Improvement +3.04 +2.43 +1.92 +1.82 +2.48 +3.78 +2.25 +4.09 +2.43 +2.24 +3.67 +4.15 +3.32
Table 1: Results on DomainNet [50], comprising $4$ domains: Real, Clipart, Sketch, and Painting, for a total of 12 domain adaptation scenarios. All models use the Swin-B [37] backbone. Best results are reported in bold, whereas second best results are underlined.
4.1 Experimental Setup
Datasets. We evaluate LAGUNA on four comprehensive UDA benchmarks: DomainNet [50], GeoPlaces [26] GeoImnet [26], and Ego2Exo [27, 20]. DomainNet consists of $400K$ images across $345$ classes and is used to evaluate performance in $12$ domain transfer settings [68, 27]. GeoImnet and GeoPlaces are subsets of the larger GeoNet dataset [26], contain over 750K images and focus on geographic disparities in image classification (GeoImNet with $600$ classes) and scene recognition (GeoPlaces with $205$ classes). Finally, Ego2Exo is a subset of EgoExo-4D curated for domain adaptation. It involves a total of $9086$ videos and $24$ action categories. DomainNet and GeoNet include per-image captions derived from metadata and BLIP-2 [32], respectively, while Ego2Exo provides video descriptions from the original EgoExo-4D dataset.
Model and training details. Our method comprises three base components: a language model for structure definition (Stage 1), a text supervisor model (Stage 2), and a visual backbone (Stage 3). We use a pre-trained SentenceTransformer [55] for structure definition and a BERT base model [59] as text supervisor. The latter is trained on the source domain of each scenario for 5 epochs with a batch size of 64, a learning rate of 1e-4, and the AdamW [41] optimizer. The visual encoders are tailored to each dataset. For DomainNet, we adapt the Swin-B backbone [37], following prior work [77, 73, 25]. For GeoImnet and GeoPLaces, we utilize the ViT-B backbone [12] following [68, 77]. For Ego2Exo, we employ pre-extracted Omnivore-L features [20] as in [27]. All vision models undergo joint training with the other network components for $10$ epochs with a batch size of $32$, an initial learning rate of $41e-$ 4, and cosine scheduling. We set the loss weights to $\lambda_{1}$=1.0 $\lambda_{2}$=0.1 and $\lambda_{3}$=0.001 to normalize the relative magnitude of the related loss functions.
Model Geolmnet GeoPlaces Avg.
[…]
Source Only 52.46 51.91 44.90 36.85 46.53
CDAN [40] 54.48 53.87 42.88 36.21 46.86
MemSAC [25] 53.02 54.37 42.05 38.33 46.94
ToAlign [68] 55.67 55.92 42.32 38.40 48.08
MDD [75] 51.57 50.73 42.54 39.23 46.02
DALN [7] 55.36 55.77 41.06 40.41 48.15
PMTrans [77] 56.76 57.60 46.18 40.33 50.22
UDA with Language Guidance
TextMatch [27] 49.68 54.82 53.06 50.11 51.92
nGramMatch [27] 49.53 51.02 51.70 49.87 50.93
LaGTran [27] 63.67 64.16 56.14 57.02 60.24
LAGUNA 67.39 69.97 61.15 63.51 65.40
Improvement +3.72 +5.81 +5.01 +6.49 +5.26
Table 2: Results on GeoImnet [26] and GeoPlaces [26] with 4 adaptation scenarios. Best results are in bold, whereas second best results are underlined. All methods use the ViT-B [12] backbone.
[…]
Source Only 8.39 15.66 12.03
TA3N [8] 6.92 27.95 17.44
TransVAE [69] 12.06 23.34 17.70
Zero-shot Video Recognition
EgoVLP [35] 5.89 19.35 12.62
LaVILA [76] 5.86 23.16 14.51
UDA with Language Guidance
TextMatch [27] 10.36 13.57 11.97
nGramMatch [27] 11.50 15.46 13.98
LaGTran [27] 12.34 30.76 21.55
LAGUNA 13.52 33.45 33.45
Improvement +1.18 +2.69 +1.94
Table 3: Results on the Ego2Exo benchmark from EgoExo-4D [20]. Best results are in bold, whereas second best results are underlined. All methods use official pre-extracted Omnivore-L features [17, 20] except EgoVLP [35] and LaVILA [76], which use Timesformer-B [3] backbone.
4.2 Results
We benchmark our model against various domain adaptation methods that have reported results on the considered datasets [77, 68, 25, 75, 7, 73]. Besides classic UDA methods, we also compare with recent approaches using language guidance [27]. Following the evaluation protocols of each benchmark, we report accuracy for DomainNet, GeoImNet, and GeoPlaces, while for Ego2Exo, we use per-class mean accuracy as reported in [27].
DomainNet. Table 1 presents the results on the DomainNet dataset, encompassing $12$ domain adaptation scenarios across four distinct domains: Real (R), Clipart (C), Sketch (S), and Painting (P). LAGUNA is compared against previous state-of-the-art methods utilizing the Swin-B backbone to ensure a fair comparison. Notably, LAGUNA surpasses all prior work in all $12$ scenarios, achieving an average accuracy improvement of $+3.32\%$.
[…]
These consistent improvements across all tested datasets, with a total of 18 scenarios, underscore the effectiveness of our approach both when considering classic UDA algorithms and recent methods exploring the use of language guidance to assist domain adaptation.
4.3 Ablation Study
In this section, we analyze the effect of the individual components of LAGUNA to provide an in-depth assessment of model performance and justify our design choices. Specifically, we investigate (a) the impact of our methodological elements, (b) the effects of different reference structures, (c) the behavior of LAGUNA with different ratios of pseudo-labels, and (d) the qualitative illustration of the alignment of relative and absolute spaces. More ablations are reported in the Supplementary Material.
Methodological Elements. In Table 4, we ablate the main methodological elements of LAGUNA on GeoImnet. We set up this ablation by removing the core elements of LAGUNA, namely the structure loss ($\mathcal{L}_{S}$ Eq. (6)), the use of domain-agnostic reference anchors $\mathcal{A}$, the learnable anchors $\mathcal{A}_{t/s}$, the cross-domain attention (CD Attn. - Eq. (9), (10)), and the regularization loss ($\mathcal{L}_{Reg}$ Eq. (12)). We then progressively add the considered elements until we reach LAGUNA (settings (1)-(5)). For all settings, we report the relative improvement (w.r.t. previous row) and the absolute one (w.r.t. first row). In setting (1), the visual classifier is trained on source and target samples using labels and pseudo-labels predicted by the text supervisor, which leads to a baseline model akin to LaGTran [27]. In setting (2), we add the reference anchors $\mathcal{A}$ and train the visual classifier as in setting (1), but we also add a loss to align visual representations with the related class anchors in absolute coordinates through cosine similarity. This yields an improvement of $+0.87$ with respect to (1), suggesting that imposing a reference structure beyond pseudo-labels is beneficial to performance but that absolute alignment is too restrictive. In setting (3), we extend (2) with a set of domain-independent learnable anchors ($\checkmark$*✓ ∗ in Table 4, indicating that the weights of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are shared) and add the structure-preserving loss $\mathcal{L}_{S}$ (Eq. 6). The introduction of learnable anchors enables the relative alignment of the visual domain with the reference language domain, leading to an improvement of $+1.32$ with respect to (2) and $+2.10$ over (1), which highlights the benefits of the proposed relative alignment over an absolute one. Considering domain-specific anchors in setting (4) allows source and target domains to focus on the individual characteristics of the respective domains, leading to a further improvement of $+1.77$ with respect to (3) and a robust $+3.87$ with respect to the baseline (1). Adding cross-domain attention in setting (5) allows to bridge the gap between target and source representations, leading to an additional improvement of $+0.44$ with respect to (4) and an overall improvement of $+4.31$ with respect to (1). We achieve the final configuration in LAGUNA, where we add the regularization loss, which prevents collapse in learning domain-specific anchors and leads to a further improvement of $+1.16$ with respect to (5) and an overall gain of $+5.47$ with respect to the baseline (1). Together, these improvements show the advantages of the careful design of LAGUNA, highlighting the benefits of learning domain-specific visual representations still aligned to a reference language structure.
Setting $\mathcal{L}_{S}$ $\mathcal{A}$ $\mathcal{A}_{t/s}$ CD Attn. $\mathcal{L}_{Reg}$ Avg. Acc. Rel. Imp. Abs. Imp.
(1) - - - - - 63.21 - -
(2) - - - - 63.99 +0.78 +0.78
(3) ✓* - - 65.22 +1.32 +2.10
(4) - - 67.08 +1.77 +3.87
(5) - 67.52 +0.44 +4.31
LAGUNA 68.68 +1.16 +5.47
Table 4: Ablation on LAGUNA’s methodological components. ’*’ means that the learnable anchors $\mathcal{A}_{t/s}$ are domain-independent (i.e., the same set of learnable anchors for target and source).
Ablation on reference structures. Fig. 3 delves into the importance of defining a semantic structure for our method (Stage 1). We ablate three models: SentenceTransformer[55], CLIP [54], and BERT [59], each with varying capabilities of extracting semantic relationships. The figure shows semantic similarity maps for $100$ randomly selected classes from GeoImnet and the average accuracy obtained by LAGUNA on the entire dataset when using the specified model. As can be seen, SentenceTransformer keeps low similarities for distinct classes (low values off-diagonal), BERT maintains high similarities for most classes but also presents some spurious correlations between distinct classes, while the behavior of CLIP is in-between. SentenceTransformer, which can better differentiate between semantic relations, leads to the best accuracy.
Figure 3: Ablation on three language models, SentenceTransformer, CLIP, and BERT, used to define the reference structure. We visualize the similarity maps of $100$ randomly selected classes from GeoImnet (green to yellow areas indicate higher similarity) and the average accuracies on all classes achieved by using each model to define the reference space. The best result is in bold, and the second best is underlined.
Ablation on the ratio of target samples. LAGUNA demonstrates strong generalization capabilities even with limited target domain data, as shown in Table 5. We progressively increase the amount of pseudo-labeled target data used for training (10%, 20%, 30%, 50%, 75%) and observe the impact on performance. Notably, LAGUNA achieves high accuracy even with only $10\%$ of the target samples and attains state-of-the-art results with just $20\%$. This efficiency underscores the benefits of our structure-driven approach, which organizes sample projections within domain-specific representation spaces while leveraging a reference geometrical structure to guide their placement.
Dataset 10% 20% 30% 50% 75% 100%
GeoImnet 61.40 63.43 64.21 66.42 66.71 68.68
GeoPlaces 55.35 57.85 58.95 60.02 60.75 62.33
Avg. 58.38 60.64 61.58 63.22 63.72 65.40
Table 5: Ablation using different target data ratios. We report average accuracy on GeoImnet and GeoPlaces. The result in Green indicates the ratio needed for LAGUNA to obtain SOTA accuracy.
Figure 4: Visualization of Source and Target (Cross-Domain) visual features of 1000 randomly selected samples from GeoImnet with t-SNE. We illustrate their position in latent space from two perspectives: left) absolute coordinates and right) relative coordinates. Notably, in absolute coordinates, the two representation spaces are quite distinct, adhering to our training strategy without direct alignment. Relative encodings w.r.t $\mathcal{A}_{s}$ and $\mathcal{A}_{t}$ are aligned.
Qualitative illustration of embedding spaces. LAGUNA uniquely aligns representation structures without enforcing alignment in absolute coordinates but rather leveraging a reference structure defined by a language model and the class labels to guide the alignment of inter-class affinities. Figure 4 provides a qualitative validation of this approach. Using t-SNE, we visualize the projection of $1000$ randomly selected samples from the GeoImnet validation set, where the source and the target domains belong to the USA and Asia, respectively. Notably, the features are well separated in absolute coordinates, reflecting their unique domain-related visual properties. However, when using relative encodings with domain-specific learnable anchors, the features blend together due to the structure-preserving loss $\mathcal{L}_{S}$. This demonstrates that semantic alignment can be achieved without forcing representations into a shared space, allowing each domain to preserve its unique characteristics while adhering to a common structure.
5 Conclusions
This work introduces LAGUNA, a novel domain adaptation approach leveraging geometrical structures of semantically equivariant spaces to guide adaptation. LAGUNA defines a reference representation space structure based on domain-agnostic class semantic similarities encoded by a language model, ensuring the organization of sample projections reflects this structure while preserving domain-specific characteristics. By conditioning the classifier to adhere to this structure, LAGUNA encourages structural similarity across domain-specific latent spaces, retaining unique features for improved classification. This is achieved through pseudo-labeling and learnable domain-specific anchors, guided by a loss function that prioritizes mimicking geometrical associations over direct representation alignment. Extensive experiments demonstrate LAGUNA’s superior performance compared to existing methods. This research highlights the importance of structural alignment and language-guided learning in domain adaptation. Future work could explore extending LAGUNA to more complex scenarios and multi-modal data.
Supplementary Material
Appendix A Introduction
In this document, we report additional ablation studies regarding the choice of language models and the structural guidance of the target domain. Moreover, we provide a more detailed explanation of the motivation behind our regularization loss $\mathcal{L}_{Reg}$ in Eq. (8). Finally, we attach a Python file, LAGUNA.py, demonstrating the implementation of LAGUNA in PyTorch. We will release the full code and configurations upon acceptance.
Appendix B Ablation on Language Models
In addition to the language model used for defining the reference structure through $\mathcal{A}$ in Stage 1, LAGUNA employs language models in Stage 2 for encoding image/video captions trained for text classification and capturing semantic structure. This model is then used to generate pseudo-labels and serve as structure guidance for the target domain in Stage 3. We ablate the choice of language model in these Stages, comparing LAGUNA’s performance when employing SentenceTransformer [55], CLIP [54], or BERT [59] on GeoImnet dataset in two settings: (1) using SentenceTransformer-defined $\mathcal{A}$ (Table 6) as it can better model semantic relationships (recall Fig. 3 from the main manuscript) and (2) defining $\mathcal{A}$ with the same language model as the one chosen for training (Table 7).
In setting (1), BERT outperforms CLIP and SentenceTransformer, motivating our choice for the language model. In setting (2), CLIP achieves the highest accuracy, but its overall performance remains lower than BERT’s in setting (1). These results not only emphasize the best model and justify our choice but also demonstrate that using different models for structure definition and training can boost performance since the choices are tailored for their specific characteristics. Specifically, the Stage 1 model is selected for its ability to model meaningful relationships between classes, while Stage 2 is intended for building better representations of captions that learn the defined structure and can also produce better pseudo-labels.
Scenario SentenceTransformer CLIP BERT
[…]
Avg. 65.87 68.11 68.68
Table 6: Ablation on three language models used in stage 2 and 3, SentenceTransformer, CLIP, and BERT with SentenceTransformer-defined $\mathcal{A}$ on GeoImnet with two domains: Asia (A) and Usa (U). The best result is in bold, and the second best is underlined.
Scenario SentenceTransformer CLIP BERT
[…]
Avg. 65.87 67.88 67.52
Table 7: Ablation on three language models used in stage 2 and 3, SentenceTransformer, CLIP, and BERT, used for Stage-2 training on GeoImnet with two domains: Asia (A) and Usa (U). In this scenario, the same pre-trained model encodes $\mathcal{A}$ (stage 1). The best result is in bold, and the second best is underlined.
Appendix C Ablation on structure guidance for target domain in Stage-3
During Stage-3 training of LAGUNA, in addition to classification training, we train our model also for structure-preserving as defined from $\mathcal{A}$ using the structural loss $\mathcal{L}_{S}$. In this particular training loss, for the source domain, we supervise the visual relative representation $\mathbf{r}^{g_{i}^{s}}=rel(g_{i}^{s},A_{s})$ with the language relative encoding of the class corresponding language anchor $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$. On the other hand, in the target domain, since we do not have a ground truth label but only the estimated pseudo label $\overline{y}_{i}^{t}$, we supervise the visual relative representation $\mathbf{r}^{g_{i}^{t}}=rel(g_{i}^{t},A_{t})$ with the language relative encoding of the sample captioning $\mathbf{r}^{z_{i}^{t}}=rel(z_{i}^{t},\mathcal{A})$ to benefit from the richer information expressed in the language encoding compared to the estimated pseudo-label. In this ablation, we investigate the benefits of using the $\mathbf{r}^{z_{i}^{t}}$ compared to relying on the pseudo-label and supervise the visual encoding with $\mathbf{r}^{\overline{y}_{i}^{t}}=rel(\mathcal{A}[y_{i}^{t}],\mathcal{A})$. Specifically, we ablate on the GeoImnet and Ego2Exo datasets and report the results in Table 8. Notably, $\mathbf{r}^{z_{i}^{t}}$ constantly gives better performances than $\mathbf{r}^{\overline{y}_{i}^{t}}$ in all scenarios motivating our choice to use the textual encoding rather than the pseudo-label corresponding anchor to create the relative representations for structure supervision in the target domain.
GeoImnet Ego2Exo
[…]
$\mathbf{r}^{z_{i}^{t}}$ $\mathbf{r}^{\overline{y}_{i}^{t}}$ $\mathbf{r}^{z_{i}^{t}}$ $\mathbf{r}^{\overline{y}_{i}^{t}}$ $\mathbf{r}^{z_{i}^{t}}$ $\mathbf{r}^{\overline{y}_{i}^{t}}$ $\mathbf{r}^{z_{i}^{t}}$ $\mathbf{r}^{\overline{y}_{i}^{t}}$
67.39 66.77 69.97 68.14 13.52 13.49 33.45 32.88
Table 8: Ablation on structure guidance representations on GeoImnet and Ego2Exo. We consider the comparison between relative representations obtained from using the language encoding $\mathbf{r}^{z_{i}^{t}}$ and from using $\mathbf{r}^{\overline{y}_{i}^{t}}$. The best result for each scenario is in bold. Note that for Ego2Exo the reported metric is mean per class accuracy.
Appendix D Regularization Loss Extended
In Section 3.4.1 of the main manuscript, we explain how the structure loss $\mathcal{L}_{S}$ is supported by a regularization loss $\mathcal{L}_{Reg}$ to avoid feature collapsing. The logic behind $\mathcal{L}_{Reg}$ is built upon two base concepts of linear algebra: Gram matrix and determinant. To better explain this logic and its relation with volume and feature collapsing, in this section, we give more information about what is a Gram matrix, what its determinant represents, how it is connected to volume, and why it serves as a collapsing indicator to motivate our approach.
Gram matrix. Given a set of vectors $v_{1},v_{2},\dots,v_{n}$ in $\mathbb{R}^{d}$, the Gram matrix $\gamma$ is defined as:
$\gamma=V^{T}V$
where $V$ is the matrix whose columns are the vectors $v_{i}$, and $\gamma$ is an $n\times n$ symmetric matrix whose elements are the pairwise inner products of the vectors:
$\gamma_{ij}=\langle v_{i},v_{j}\rangle$
Additionally, its rank corresponds to the number of linearly independent vectors, and its determinant (if full-rank) relates to the volume of the parallelepiped spanned by the vectors. If $\gamma$ is not full rank, the vectors used to construct it are linearly dependent, so they do not span the full n-dimensional space. Geometrically, this implies that the volume of the parallelepiped they define is zero, as they lie in a lower-dimensional subspace. In other words, a lower rank or a volume of zero of the vectors composing $\gamma$ indicates collapsing in a lower subspace. In LAGUNA, we use these properties of the Gram matrix to control and prevent the collapsing of the anchor vectors composing $\mathcal{A}_{s}$ and $\mathcal{A}_{t}$. Particularly, from Eq. (7) in the manuscript, we calculate $\gamma_{s}=A_{s}^{T}A_{s}$ and $\gamma_{s}=A_{t}^{T}A_{t}$.
Determinant of Gram matrix. To determine if a set of vectors has collapsed, one can either compute the rank of its Gram matrix or evaluate its volume in space. In LAGUNA, we choose the latter approach as we can compare it with a reference volume that helps prevent collapse during training. Specifically, to calculate the volume of the parallelepiped formed by the set of vectors (anchors) in $\mathcal{A}_{s/t}$, we utilize the determinant. The determinant of a Gram matrix represents the square of the volume of the parallelepiped defined by the vectors used to compute the matrix (e.g., $Det(\gamma_{s})$ gives the square of the volume occupied by all anchors in $\mathcal{A}_{s}$). Thus, the volume for $\mathcal{A}_{s/t}$ can be expressed as:
$\text{Volume}_{s/t}=\sqrt{Det(\gamma_{s/t})}.$
Importantly, if the anchors in $\mathcal{A}_{s/t}$ are linearly dependent (i.e., not full rank), $Det(\gamma_{s/t})=0$. To prevent collapse, we can regularize $\mathcal{L}_{S}$ by requiring the volume to remain greater than zero. However, in higher-dimensional spaces, volumes can have very large values. For numerical stability, we replace $Det(\gamma_{s/t})$ with $logDet(\gamma_{s/t})$ in Eq. (8) and enforce that the space occupied by anchors in $\mathcal{A}_{s/t}$ remains log comparable to that of the reference anchors in $\mathcal{A}$ preventing their collapse into a zero volume. Specifically, Eq. (8) in the manuscript is defined as follows:
$\mathcal{L}_{Reg}$ $=$ $|logDet(\gamma_{t})-logDet(\gamma)|+$ (12)
$|logDet(\gamma_{s})-logDet(\gamma)|,$
where $\gamma=\mathcal{A}^{T}\mathcal{A}$.

### v3

@@
 Unsupervised domain adaptation remains a critical challenge in enabling the knowledge transfer of models across unseen domains.
-Existing methods struggle to balance the need for domain-invariant representations with preserving domain-specific features, which is often due to alignment approaches that impose the projection of samples with similar semantics close in the latent space despite their drastic domain differences.
+Existing methods struggle to balance the need for domain-invariant representations with preserving domain-specific features.
+This is often due to the fact that alignment approaches impose the projection of samples with similar semantics close to each other in the latent space, despite drastic domain differences.
 We introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach that shifts the focus from aligning representations in absolute coordinates to aligning the relative positioning of equivalent concepts in latent spaces.
-LAGUNA defines a domain-agnostic structure upon the semantic/geometric relationships between class labels in language space and guides adaptation, ensuring that the organization of samples in visual space reflects reference inter-class relationships while preserving domain-specific characteristics.
-Remarkably, LAGUNA surpasses previous works in $18$ different adaptation scenarios across four diverse image and video datasets with average accuracy improvements of $+3.32\%$ on DomainNet, $+5.75\%$ in GeoPlaces, $+4.77\%$ on GeoImnet, and $+1.94\%$ mean class accuracy improvement on EgoExo4D.
+LAGUNA defines a domain-agnostic structure on the semantic/geometric relationships between class labels in the language space and guides the adaptation, ensuring that the organization of samples in visual space reflects reference inter-class relationships while preserving domain-specific characteristics.
+Remarkably, LAGUNA surpasses previous work in $18$ different adaptation scenarios across four diverse image and video datasets with average accuracy improvements of up to $+3.32\%$ on DomainNet, $+5.75\%$ on GeoPlaces, $+4.77\%$ on GeoImnet, and $+1.94\%$ mean class accuracy improvement on EgoExo4D.
 1 Introduction
@@
 This assumes the availability for both source and target samples of natural language descriptions, which can be generated from available captioning models [33] or collected from the web at a fraction of the cost of typical human labeling procedures [27].
-Specifically, LAGUNA employs a 3-stage approach to structurally align the source and target representation spaces while allowing each domain to preserve its typical patterns.
+LAGUNA employs a 3-stage approach to structurally align source and target representation spaces while allowing domains to preserve typical patterns (Figure 2).
 In Stage 1, textual class labels are mapped to a domain-agnostic reference space representing their relationships at the conceptual level.
-In Stage 2, a language model is trained to map textual captions to the reference latent space in order to provide pseudo-labels for target samples Finally, in Stage 3, a cross-domain classifier is trained to encourage domain-specific representations to follow the reference structure.
-Experiments demonstrate the superiority of LAGUNA over existing state-of-the-art methods across $18$ different domain adaptation scenarios sourced from four diverse image and video datasets, with gains of $+3.32\%$ on DomanNet [50], $+5.75\%$ on GeoPlaces [26], $+4.77\%$ on GeoImnet [26] and $+1.94\%$ mean per class accuracy on EgoExo4D [20].
+In Stage 2, a language model is trained to map textual captions to the reference latent space, providing pseudo-labels for target samples.
+In Stage 3, a cross-domain classifier is trained to encourage domain-specific representations to follow the reference structure.
+Experiments demonstrate the superiority of LAGUNA over existing state-of-the-art methods across $18$ different domain adaptation scenarios sourced from four diverse image and video datasets, with gains of up to $+3.32\%$ on DomanNet [52], $+5.75\%$ on GeoPlaces [26], $+4.77\%$ on GeoImnet [26] and $+1.94\%$ mean per class accuracy on EgoExo4D [20].
 We further report ablations to analyze the specific contributions of our design choices.
 In sum, our main contributions are as follows: 1) we investigate the use of relative representations for UDA, showing its advantages with respect to the current absolute alignment paradigm; 2) we propose LAGUNA, a three-stage method which learns a cross-domain classifier where individual source and target spaces are distinct yet aligned to a reference structure derived from language; 3) through extensive ablations and comparisons with state-of-the-art models, we show the superiority of the proposed approach.
-Code available at: https://github.com/ADiko1997/LAGUNA.git
 2 Related Work
@@
 The model is trained to be structurally aligned to anchors ${}^{*}\mathcal{A}$ with a structural loss $\mathcal{L}_{S}$ and to predict class labels with a cross-entropy loss $\mathcal{L}_{CE}$.
-In Stage 3, a cross-domain visual classifier is trained on labeled target and unlabeled source images.
-A visual encoder $\mathcal{V}$ is used to represent target and source samples as $g_{i}^{t}$ and $g_{i}^{s}$.
-The cosine similarities $\mathbf{r}^{g_{i}^{t}}$ between the target visual representation $g_{i}^{t}$ and all learnable target anchors $\mathcal{A}_{t}$ are bound to be similar to the cosine similarities $\mathbf{r}^{z_{i}^{t}}$ between the textual representation of the target caption $z_{i}^{t}$ predicted through the language model $\mathcal{G}$ and the reference domain-agnostic anchors ${}^{*}\mathcal{A}$ by means of the structural loss $\mathcal{L}_{S}$.
-A similar processing is applied to the visual representation of source sample $g_{i}^{s}$, the textual representation of the ground truth class ${}^{*}A[y_{i}^{S}]$ and the learnable source anchors $\mathcal{A}_{s}$.
+In Stage 3, a cross-domain visual classifier is trained on labeled source and unlabeled target images.
+A visual encoder $\mathcal{V}$ is used to represent target and source samples as $g_{i}^{t}$ and $g_{i}^{s}$, respectively.
+The cosine similarities $\mathbf{r}^{g_{i}^{t}}$ between $g_{i}^{t}$ and learnable target anchors $\mathcal{A}_{t}$ are bound to be alike to the similarities $\mathbf{r}^{z_{i}^{t}}$ between textual representation of the target caption $z_{i}^{t}$ predicted via $\mathcal{G}$ and ${}^{*}\mathcal{A}$ by means of $\mathcal{L}_{S}$.
+A similar processing is applied to $g_{i}^{s}$, the textual representation of the ground truth class ${}^{*}A[y_{i}^{S}]$ and $\mathcal{A}_{s}$.
 A regularization loss $\mathcal{L}_{Reg}$ is used to avoid collapse during anchor learning.
-Representations $g_{i}^{t}$ and $g_{i}^{s}$ are hence grounded into the learnable source anchors $\mathcal{A}_{s}$ through a novel Cross-Domain Attention layer taking visual representations as queries, source/target anchors as keys and source anchors as values.
-Classification is finally achieved through an MLP classifier supervised through a cross-entropy loss $\mathcal{L}_{CE}$ using ground truth labels in the source domain and pseudo-labels in the target domain.
-The visual encoder, Cross-Domain Attention, and MLP layers share weights across domains.
+Representations $g_{i}^{t}$ and $g_{i}^{s}$ are hence grounded into the learnable $\mathcal{A}_{s}$ through a novel Cross-Domain Attention layer taking visual representations as queries, source/target anchors as keys and source anchors as values.
+Classification is finally achieved with an MLP classifier supervised via $\mathcal{L}_{CE}$ using ground truth labels in the source domain and pseudo-labels in the target domain.
 2.2 Language Guidance in Vision UDA
@@
 3 Method
+LAGUNA comprises a 3-stage process, as illustrated in Fig.
+2.
 In Stage 1, a language-based reference structure space is defined.
 We represent this structure as a set of anchors, which we will use as a reference to guide representation learning.
-In Stage 2, a text-supervisor language model is trained to map image/video captions to class categories in order to provide pseudo-labels for the unlabeled target domain.
-The language model is also trained to keep the internal text embeddings aligned to the domain-agnostic structure defined in Stage 1.
-Finally, in Stage 3, a cross-domain visual classifier is trained to categorize source and target samples by learning domain-specific action anchors bound to be structurally similar to the domain-agnostic structure defined in Stage 1.
+In Stage 2, a text-supervisor language model is trained to map image/video captions to class categories, providing pseudo-labels for the unlabeled target domain.
+The language model is also trained to keep internal text embeddings aligned to the domain-agnostic structure defined in Stage 1.
+In Stage 3, a cross-domain visual classifier is trained to categorize source and target samples by learning domain-specific action anchors that are structurally similar to the domain-agnostic structure defined in Stage 1.
 3.1 Problem Setup
@@
 $z_{i}^{s}=\mathcal{G}(l_{i}^{s}).$ (1)
-Next, we encourage the vector representation $z_{i}^{s}$ to be geometrically aligned to the anchor $\mathcal{A}[y_{i}^{s}]$ corresponding to ground truth action $y_{i}^{s}$.
+Next, we encourage the vector representation $z_{i}^{s}$ to be geometrically aligned to $\mathcal{A}[y_{i}^{s}]$ corresponding to ground truth action $y_{i}^{s}$.
 To do so, we compute $\mathbf{r}^{z_{i}^{s}}=rel(z_{i}^{s},\mathcal{A})$, the relative encoding of $z_{i}^{s}$, and supervise it to be similar to $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$, the relative encoding of the anchor $\mathcal{A}[y_{i}^{s}]$, with the following structure-preserving loss:
@@
 3.4 Stage 3: Cross-Domain Visual Classifier
-Stage 3 aims to train a cross-domain visual classifier aligning representations extracted through a visual encoder $\mathcal{V}$ to the structure imposed by the domain-agnostic language anchors $\mathcal{A}$.
+Stage 3 trains a cross-domain visual classifier aligning representations extracted through a visual encoder $\mathcal{V}$ to the structure imposed by the domain-agnostic language anchors $\mathcal{A}$.
 We allow each domain to develop its own latent space, but we encourage both spaces to be aligned to two sets of learnable anchors $\mathcal{A}_{t}\in\mathbb{R}^{N_{c}\times D_{v}}$ and $\mathcal{A}_{s}\in\mathbb{R}^{N_{c}\times D_{v}}$ specific to the target and source domain respectively.
-We further ensure that the structures of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are aligned to the domain-agnostic, fixed anchors $\mathcal{A}$ through the supervision provided by ground truth labels $y_{i}^{s}$ in the source domain and the text supervision provided by $\mathcal{G}$ in the target domain in the form of the predicted vector $z_{i}^{t}$ and the predicted pseudo-label $\overline{y}_{i}^{t}$.
+We further ensure that the structures of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are aligned to the structure of $\mathcal{A}$ through the supervision provided by ground truth labels $y_{i}^{s}$ in the source domain and the text supervision provided by $\mathcal{G}$ in the target domain in the form of the predicted vector $z_{i}^{t}$ and the predicted pseudo-label $\overline{y}_{i}^{t}$.
 The source and target datasets are merged ($\hat{\mathcal{X}}=\mathcal{X}_{t}+\mathcal{X}_{s}$) and used for training with a total of $M=N_{s}+N_{t}$ samples.
@@
 Target and source images are processed by $\mathcal{V}$ to obtain latent representations $g_{i}^{t}$ and $g_{i}^{s}$, while target captions $l_{i}^{t}$ are processed by $\mathcal{G}$ to obtain latent representations $z_{i}^{t}$ as follows:
-$g_{i}^{t}=\mathcal{V}(x_{i}^{t}),\hfill z_{i}^{t}=\mathcal{G}(l_{i}^{t}),\hfill g_{i}^{s}=\mathcal{V}(x_{i}^{s}).$ (5)
+$g_{i}^{t}=\mathcal{V}(x_{i}^{t}),\hskip 14.22636ptz_{i}^{t}=\mathcal{G}(l_{i}^{t}),\hskip 14.22636ptg_{i}^{s}=\mathcal{V}(x_{i}^{s}).$ (5)
 These representations and the anchors are used to achieve two main training objectives: 1) structure learning and 2) classification training.
@@
 $\mathcal{L}_{S}=\begin{cases}|\mathbf{r}^{g_{i}^{t}}-\mathbf{r}^{z_{i}^{t}}|,&\textit{if Target Domain}\\ |\mathbf{r}^{g_{i}^{s}}-\mathbf{r}^{y_{i}^{s}}|,&\textit{otherwise}\end{cases}.$ (6)
-This loss encourages the visual representations to maintain relationships with their respective domain-specific visual anchors that mirror the relationships expressed by the domain-invariant anchors $\mathcal{A}$.
+This loss encourages visual representations to maintain relationships with their respective domain-specific visual anchors that mirror the relationships expressed by the domain-invariant anchors $\mathcal{A}$.
 This process supervises both the learnable anchors $\mathcal{A}_{s},\mathcal{A}_{t}$ and the visual encoder $\mathcal{V}$.
 However, since this loss alone might lead to the collapse of anchor representations into a small region, hindering effective classification, we introduce a volume (or spread) regularization loss.
-This loss treats each set of anchors as a multidimensional parallelotope whose spread or volume occupied in the latent space can be measured by the determinant of its Gram matrix.
+This loss treats each set of anchors as a multidimensional parallelotope whose volume in the latent space can be measured by the determinant of its Gram matrix.
 Given the Gram matrices for the three sets of anchors:
-$\gamma=\mathcal{A}\mathcal{A}^{T},\hfill\gamma_{s}=\mathcal{A}_{s}\mathcal{A}_{s}^{T},\hfill\gamma_{t}=\mathcal{A}_{t}\mathcal{A}_{t}^{T},$ (7)
-we encourage the volume of each domain-specific parallelotope to be approximately equal to the volume of the reference anchors $\mathcal{A}$ with the following regularization loss: e
+$\gamma=\mathcal{A}\mathcal{A}^{T},\hskip 14.22636pt\gamma_{s}=\mathcal{A}_{s}\mathcal{A}_{s}^{T},\hskip 14.22636pt\gamma_{t}=\mathcal{A}_{t}\mathcal{A}_{t}^{T},$ (7)
+we encourage the volume of each domain-specific parallelotope to be approximately equal to the volume of the reference anchors $\mathcal{A}$ with the following regularization loss:
 $\mathcal{L}_{Reg}$ $=$ $|logDet(\gamma_{t})-logDet(\gamma)|+$ (8)
 $|logDet(\gamma_{s})-logDet(\gamma)|,$
-where we consider log-determinants instead of determinants for numerical stability.
-This regularization loss ensures the visual anchors occupy a volume similar to the one of $\mathcal{A}$, preventing collapse and promoting better separation of classes in the representation space.
+where we use log-determinants for numerical stability.
+This regularization loss ensures taht visual anchors occupy a volume similar to the one of $\mathcal{A}$, preventing collapse and promoting better class separation in the representation space.
 3.4.2 Classification Training
-The network undergoes classification training concurrently with structure learning.
+We perform classification training concurrently with structure learning.
 To ensure that source and target visual representations are grounded in the structure imposed by $\mathcal{A}$, while still being free to capture domain-specific nuances, we propose a novel cross-domain attention layer that aims to ground domain-specific visual representations $g_{i}^{t}$ and $g_{i}^{s}$ to the structure of source anchors $\mathcal{A}_{s}$.
-We include two Cross-Domain attention layers sharing the same weights for the target and source branches of the architecture (See Fig.
+We include two Cross-Domain attention layers sharing the same weights for the target and source branches of LAGUNA (See Fig.
 2).
@@
 This process results in a unified representation since the output of the attention layer always depends on values coming from the source anchors $\mathcal{A}_{s}$.
-The residual connection included in the attention layer is introduced to account for domain-dependent class characteristics, which results in final feature vectors $f_{i}^{s}$ and $f_{i}^{t}$.
+The residual connection included in the attention layer is introduced to account for domain-dependent class characteristics, which results in the feature vectors $f_{i}^{s}$ and $f_{i}^{t}$.
 Finally, an MLP classification head processes $f_{i}^{t}$ and $f_{i}^{s}$ to output predictions $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{s}$.
 The MLP has shared weights across domains and is optimized with a cross-entropy loss using the ground truth label $y_{i}^{s}$ for source examples and the pseudo-label $\overline{y}_{i}^{t}$ predicted by $\mathcal{G}$ in the case of target samples.
-The overall training objective of the classifier is the following:
-$\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(\hat{y_{i}},y_{i}^{*})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{*},\mathbf{r}^{g_{i}})+\lambda_{3}\mathcal{L}_{Reg}(\gamma,\gamma_{*}),$ (11)
+The overall training objective is:
+$\small\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(\hat{y_{i}},y_{i}^{*})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{*},\mathbf{r}^{g_{i}})+\lambda_{3}\mathcal{L}_{Reg}(\gamma,\gamma_{*}),$ (11)
 where $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ are hyperparameter weights to calibrate the magnitude of each loss value, and ’*’ signifies that the argument is domain-dependent [^2] For example, for $\mathcal{L}_{CE}$ the ground truth can either be $y_{i}^{s}$ or $\overline{y}_{i}^{t}$..
@@
 […]
+Swin-B Language-free Approaches
 Source Only 63.02 49.47 60.48 70.52 56.09 52.53 70.42 65.91 54.47 73.34 60.09 48.25 60.38
-MCD [58] 39.40 25.20 41.20 44.60 31.20 25.50 34.50 37.30 27.20 48.10 31.10 22.80 34.01
 MDD [77] 52.80 41.20 47.80 52.50 42.10 40.70 54.20 54.30 43.10 51.20 43.70 41.70 47.11
-CGDM [13] 49.40 38.20 47.20 53.50 36.90 35.30 55.60 50.10 43.70 59.40 37.70 33.50 45.04
 SCDA [34] 54.00 42.50 51.90 55.00 44.10 39.30 53.20 55.60 44.70 56.20 44.10 42.00 48.55
@@
 Improvement +3.04 +2.43 +1.92 +1.82 +2.48 +3.78 +2.25 +4.09 +2.43 +2.24 +3.67 +4.15 +3.32
-Table 1: Results on DomainNet [50], comprising $4$ domains: Real, Clipart, Sketch, and Painting, for a total of 12 domain adaptation scenarios.
-All models use the Swin-B [37] backbone.
-Best results are reported in bold, whereas second best results are underlined.
+ViT-B CLIP* [56] 72.39 60.90 66.81 81.37 60.90 66.81 81.37 72.39 66.81 81.37 72.39 69.90 70.38
+DAPrompt* [16] 73.90 65.90 70.40 84.90 65.80 70.20 84.60 73.50 69.90 84.90 73.80 65.80 73.80
+PADCLIP [31] 76.40 67.50 72.70 84.20 68.10 71.10 83.60 76.30 71.70 83.50 75.40 67.20 74.81
+[…]
+UniMoS* [36] 77.50 68.00 72.50 86.00 68.50 72.30 85.90 77.80 72.60 85.80 77.20 68.20 76.03
+LAGUNA 79.53 73.18 74.24 86.04 73.54 74.98 85.81 80.24 74.43 86.57 79.89 74.11 78.54
+Improvement +2.03 +4.07 +1.74 +0.04 +5.03 +2.68 -0.09 +2.44 +1.83 +0.77 +2.69 +5.91 +2.51
+Table 1: Results on DomainNet [52] in $4$ domains: Real, Clipart, Sketch, and Painting, for 12 adaptation scenarios using the Swin-B [39] and ViT-B [12] backbones.
+The best results are in bold, second bests are underlined.‘*‘ method uses domain-aware prompts, $\dagger$ our run.
 4.1 Experimental Setup
@@
 GeoImnet and GeoPlaces are subsets of the larger GeoNet dataset [26], contain over 750K images and focus on geographic disparities in image classification (GeoImNet with $600$ classes) and scene recognition (GeoPlaces with $205$ classes).
-Finally, Ego2Exo is a subset of EgoExo-4D curated for domain adaptation.
+Finally, Ego2Exo is a subset of EgoExo-4D [20] curated for domain adaptation.
 It involves a total of $9086$ videos and $24$ action categories.
-DomainNet and GeoNet include per-image captions derived from metadata and BLIP-2 [32], respectively, while Ego2Exo provides video descriptions from the original EgoExo-4D dataset.
+DomainNet and GeoNet include per-image captions derived from BLIP-2 [33] and web metadata, respectively, while Ego2Exo provides video descriptions from the original EgoExo-4D dataset.
 Model and training details.
-Our method comprises three base components: a language model for structure definition (Stage 1), a text supervisor model (Stage 2), and a visual backbone (Stage 3).
-We use a pre-trained SentenceTransformer [55] for structure definition and a BERT base model [59] as text supervisor.
-The latter is trained on the source domain of each scenario for 5 epochs with a batch size of 64, a learning rate of 1e-4, and the AdamW [41] optimizer.
-The visual encoders are tailored to each dataset.
-For DomainNet, we adapt the Swin-B backbone [37], following prior work [77, 73, 25].
+We use a pre-trained SentenceTransformer [57] for structure definition (Stage 1) and a BERT base model [61] as text supervisor (Stage 2).
+The latter is trained on the source domain of each scenario for 5 epochs with a batch size of 64, a lr of 1e-4, and the AdamW [43] optimizer.
+The visual encoders (Stage 3) are tailored to each dataset for fair comparisons with existing approaches.
+For DomainNet, we adapt the Swin-B and ViT-B backbones [39, 12].
 For GeoImnet and GeoPLaces, we utilize the ViT-B backbone [12] following [70, 79].
@@
 Table 2: Results on GeoImnet [26] and GeoPlaces [26] with 4 adaptation scenarios. Best results are in bold, whereas second best results are underlined. All methods use the ViT-B [12] backbone.
+4.2 Results
+We benchmark our model against various language-free domain adaptation methods that have reported results on the considered datasets [79, 70, 25, 77, 7, 75].
+Besides classic UDA methods, we also compare with previous approaches using language guidance [27, 56, 16, 31, 36].
+Following the evaluation protocols of each benchmark, we report accuracy for DomainNet, GeoImNet, and GeoPlaces, and per-class mean accuracy for Ego2Exo.
+DomainNet.
+Table 1 presents the results on the DomainNet dataset, encompassing $12$ domain adaptation scenarios across four distinct domains: Real (R), Clipart (C), Sketch (S), and Painting (P).
+LAGUNA is compared against previous state-of-the-art methods utilizing the Swin-B and ViT-B backbones as originally adopted by the compared methods to ensure fair comparison.
+LAGUNA significantly outperforms previous methods in 12 scenarios, with an average performance increase of $3.32\%$ for the Swin-B backbone.
+For ViT-B, LAGUNA leads in 11 of 12 scenarios, boosting performance by an average of $2.51\%$.
+Importantly, LAGUNA achieves SOTA ViT-B performance without relying on extensive CLIP pretraining and domain-specific prompting, unlike previous models [31, 36, 16].
+[…]
+[…]
+These consistent improvements across all tested datasets, with a total of 18 scenarios, underscore the effectiveness of LAGUNA with respect to both language-free and language-guided UDA approaches.
 […]
@@
 LaGTran [27] 12.34 30.76 21.55
-LAGUNA 13.52 33.45 33.45
+LAGUNA 13.52 33.45 23.49
+Target Supervised (oracle) 18.08 35.12 26.60
 Improvement +1.18 +2.69 +1.94
@@
 All methods use official pre-extracted Omnivore-L features [17, 20] except EgoVLP [37] and LaVILA [78], which use Timesformer-B [3] backbone.
-4.2 Results
-We benchmark our model against various domain adaptation methods that have reported results on the considered datasets [77, 68, 25, 75, 7, 73].
-Besides classic UDA methods, we also compare with recent approaches using language guidance [27].
-Following the evaluation protocols of each benchmark, we report accuracy for DomainNet, GeoImNet, and GeoPlaces, while for Ego2Exo, we use per-class mean accuracy as reported in [27].
-DomainNet.
-Table 1 presents the results on the DomainNet dataset, encompassing $12$ domain adaptation scenarios across four distinct domains: Real (R), Clipart (C), Sketch (S), and Painting (P).
-LAGUNA is compared against previous state-of-the-art methods utilizing the Swin-B backbone to ensure a fair comparison.
-Notably, LAGUNA surpasses all prior work in all $12$ scenarios, achieving an average accuracy improvement of $+3.32\%$.
-[…]
-[…]
-These consistent improvements across all tested datasets, with a total of 18 scenarios, underscore the effectiveness of our approach both when considering classic UDA algorithms and recent methods exploring the use of language guidance to assist domain adaptation.
 4.3 Ablation Study
-In this section, we analyze the effect of the individual components of LAGUNA to provide an in-depth assessment of model performance and justify our design choices.
-Specifically, we investigate (a) the impact of our methodological elements, (b) the effects of different reference structures, (c) the behavior of LAGUNA with different ratios of pseudo-labels, and (d) the qualitative illustration of the alignment of relative and absolute spaces.
+We analyze the effect of the individual components of LAGUNA to provide an in-depth assessment of model performance and justify design choices.
+We investigate (a) the impact of our methodological elements, (b) the effects of different reference structures, (c) the behavior of LAGUNA with different ratios of pseudo-labels, and (d) the qualitative illustration of the alignment of relative and absolute spaces.
 More ablations are reported in the Supplementary Material.
-Methodological Elements.
-In Table 4, we ablate the main methodological elements of LAGUNA on GeoImnet.
-We set up this ablation by removing the core elements of LAGUNA, namely the structure loss ($\mathcal{L}_{S}$ Eq.
-(6)), the use of domain-agnostic reference anchors $\mathcal{A}$, the learnable anchors $\mathcal{A}_{t/s}$, the cross-domain attention (CD Attn.
-- Eq.
-(9), (10)), and the regularization loss ($\mathcal{L}_{Reg}$ Eq.
-(12)).
-We then progressively add the considered elements until we reach LAGUNA (settings (1)-(5)).
-For all settings, we report the relative improvement (w.r.t.
-previous row) and the absolute one (w.r.t.
-first row).
-In setting (1), the visual classifier is trained on source and target samples using labels and pseudo-labels predicted by the text supervisor, which leads to a baseline model akin to LaGTran [27].
-In setting (2), we add the reference anchors $\mathcal{A}$ and train the visual classifier as in setting (1), but we also add a loss to align visual representations with the related class anchors in absolute coordinates through cosine similarity.
-This yields an improvement of $+0.87$ with respect to (1), suggesting that imposing a reference structure beyond pseudo-labels is beneficial to performance but that absolute alignment is too restrictive.
-In setting (3), we extend (2) with a set of domain-independent learnable anchors ($\checkmark$*✓ ∗ in Table 4, indicating that the weights of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are shared) and add the structure-preserving loss $\mathcal{L}_{S}$ (Eq.
-6).
-The introduction of learnable anchors enables the relative alignment of the visual domain with the reference language domain, leading to an improvement of $+1.32$ with respect to (2) and $+2.10$ over (1), which highlights the benefits of the proposed relative alignment over an absolute one.
-Considering domain-specific anchors in setting (4) allows source and target domains to focus on the individual characteristics of the respective domains, leading to a further improvement of $+1.77$ with respect to (3) and a robust $+3.87$ with respect to the baseline (1).
-Adding cross-domain attention in setting (5) allows to bridge the gap between target and source representations, leading to an additional improvement of $+0.44$ with respect to (4) and an overall improvement of $+4.31$ with respect to (1).
-We achieve the final configuration in LAGUNA, where we add the regularization loss, which prevents collapse in learning domain-specific anchors and leads to a further improvement of $+1.16$ with respect to (5) and an overall gain of $+5.47$ with respect to the baseline (1).
-Together, these improvements show the advantages of the careful design of LAGUNA, highlighting the benefits of learning domain-specific visual representations still aligned to a reference language structure.
 Setting $\mathcal{L}_{S}$ $\mathcal{A}$ $\mathcal{A}_{t/s}$ CD Attn. $\mathcal{L}_{Reg}$ Avg. Acc. Rel. Imp. Abs. Imp.
@@
 ’*’ means that the learnable anchors $\mathcal{A}_{t/s}$ are domain-independent (i.e., the same set of learnable anchors for target and source).
-Ablation on reference structures.
+Methodological Elements.
+Table 4, ablates the main methodological elements of LAGUNA on GeoImnet.
+We set up this ablation by removing LAGUNA’s core elements, namely the structure loss ($\mathcal{L}_{S}$ Eq.
+(6)), domain-agnostic reference anchors $\mathcal{A}$, learnable anchors $\mathcal{A}_{t/s}$, cross-domain attention (CD Attn.
+- Eq.
+(9), (10)), and the regularization loss ($\mathcal{L}_{Reg}$ Eq.
+(12)).
+We then progressively add these elements (settings (1)-(5)) until we reach LAGUNA.
+For all settings, we report the relative improvement (w.r.t.
+previous row) and the absolute one (w.r.t.
+first row).
+Absolute alignment In setting (1), the visual classifier is trained on source and target samples using labels and pseudo-labels predicted by the text supervisor, which leads to a baseline model akin to LaGTran [27].
+In setting (2), we add the reference anchors $\mathcal{A}$ and train the visual classifier as in setting (1), but we also add a loss to align visual representations with the related class anchors in absolute coordinates through cosine similarity.
+This yields an improvement of $+0.87$ with respect to (1), suggesting that imposing a reference structure beyond pseudo-labels is beneficial to performance and that absolute alignment is too restrictive.
+Relative alignment In setting (3), we extend (2) with a set of domain-independent learnable anchors ($\checkmark$*✓ ∗ in Table 4, indicating that the weights of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are shared) and add the structure-preserving loss $\mathcal{L}_{S}$ (Eq.
+6).
+The introduction of learnable anchors enables the relative alignment of the visual domain with the reference language domain, leading to an improvement of $+1.32$ with respect to (2) and $+2.10$ over (1), which highlights the benefits of the proposed relative alignment over an absolute one.
+Domain-Specific Anchors Adding domain-specific anchors in setting (4) allows source and target domains to focus on the individual characteristics of the respective domains, leading to a further improvement of $+1.77$ with respect to (3) and a robust $+3.87$ with respect to the baseline (1).
+Cross-Domain Attention Adding cross-domain attention in setting (5) bridges the gap between target and source representations, leading to a of $+0.44$ with respect to (4) and an overall improvement of $+4.31$ with respect to (1).
+Regularization Loss We achieve the final configuration in LAGUNA, where we add the regularization loss, which prevents collapse in learning domain-specific anchors and leads to a further improvement of $+1.16$ with respect to (5) and an overall gain of $+5.47$ with respect to (1).
+Together, these improvements show the effectiveness of LAGUNA, highlighting the benefits of learning domain-specific visual representations aligned to a reference language structure.
+Language models for reference structures.
 Fig.
-3 delves into the importance of defining a semantic structure for our method (Stage 1).
-We ablate three models: SentenceTransformer[55], CLIP [54], and BERT [59], each with varying capabilities of extracting semantic relationships.
-The figure shows semantic similarity maps for $100$ randomly selected classes from GeoImnet and the average accuracy obtained by LAGUNA on the entire dataset when using the specified model.
-As can be seen, SentenceTransformer keeps low similarities for distinct classes (low values off-diagonal), BERT maintains high similarities for most classes but also presents some spurious correlations between distinct classes, while the behavior of CLIP is in-between.
-SentenceTransformer, which can better differentiate between semantic relations, leads to the best accuracy.
-Figure 3: Ablation on three language models, SentenceTransformer, CLIP, and BERT, used to define the reference structure.
-We visualize the similarity maps of $100$ randomly selected classes from GeoImnet (green to yellow areas indicate higher similarity) and the average accuracies on all classes achieved by using each model to define the reference space.
-The best result is in bold, and the second best is underlined.
-Ablation on the ratio of target samples.
-LAGUNA demonstrates strong generalization capabilities even with limited target domain data, as shown in Table 5.
-We progressively increase the amount of pseudo-labeled target data used for training (10%, 20%, 30%, 50%, 75%) and observe the impact on performance.
-Notably, LAGUNA achieves high accuracy even with only $10\%$ of the target samples and attains state-of-the-art results with just $20\%$.
-This efficiency underscores the benefits of our structure-driven approach, which organizes sample projections within domain-specific representation spaces while leveraging a reference geometrical structure to guide their placement.
-Dataset 10% 20% 30% 50% 75% 100%
-GeoImnet 61.40 63.43 64.21 66.42 66.71 68.68
-GeoPlaces 55.35 57.85 58.95 60.02 60.75 62.33
-Avg. 58.38 60.64 61.58 63.22 63.72 65.40
-Table 5: Ablation using different target data ratios. We report average accuracy on GeoImnet and GeoPlaces. The result in Green indicates the ratio needed for LAGUNA to obtain SOTA accuracy.
-Figure 4: Visualization of Source and Target (Cross-Domain) visual features of 1000 randomly selected samples from GeoImnet with t-SNE.
-We illustrate their position in latent space from two perspectives: left) absolute coordinates and right) relative coordinates.
-Notably, in absolute coordinates, the two representation spaces are quite distinct, adhering to our training strategy without direct alignment.
-Relative encodings w.r.t $\mathcal{A}_{s}$ and $\mathcal{A}_{t}$ are aligned.
-Qualitative illustration of embedding spaces.
-LAGUNA uniquely aligns representation structures without enforcing alignment in absolute coordinates but rather leveraging a reference structure defined by a language model and the class labels to guide the alignment of inter-class affinities.
-Figure 4 provides a qualitative validation of this approach.
-Using t-SNE, we visualize the projection of $1000$ randomly selected samples from the GeoImnet validation set, where the source and the target domains belong to the USA and Asia, respectively.
-Notably, the features are well separated in absolute coordinates, reflecting their unique domain-related visual properties.
-However, when using relative encodings with domain-specific learnable anchors, the features blend together due to the structure-preserving loss $\mathcal{L}_{S}$.
-This demonstrates that semantic alignment can be achieved without forcing representations into a shared space, allowing each domain to preserve its unique characteristics while adhering to a common structure.
+3 reports semantic similarity maps for $100$ randomly selected classes form GeoImnet when three different language models are used in stage 1 (SentenceTransformer [57], CLIP [56], and BERT [61]), together with the average accuracy obtained keeping stages 2-3 fixed.
+Notably, accuracy follows the ability of models to distinguish distinct classes, with low off-diagonal values indicating more discriminative models.
+While we achieve best results with SentenceTransformer, LAGUNA’s performance is robust to different reference structures, maintaining SOTA results also when the same language model is used across all stages (i.e., BERT).
+Figure 3: Similarity maps of $100$ randomly selected classes from GeoImnet (yellow for high similarity) and average accuracies.
+Ratio of target samples.
+LAGUNA demonstrates strong generalization capabilities even with limited target domain data (Table 5).
+We progressively increase the number of randomly selected pseudo-labeled target data used for training (10%-75%) and observe the impact on performance.
+Notably, LAGUNA achieves high accuracy even with only $10\%$ of the samples and attains SOTA results with just $20\%$ (compared with  [27]).
+Dataset 10% 20% 30% 50% 75% 100% [27]
+GeoImnet 61.40 63.43 64.21 66.42 66.71 68.68 63.92
+GeoPlaces 55.35 57.85 58.95 60.02 60.75 62.33 56.58
+Avg. 58.38 60.64 61.58 63.22 63.72 65.40 60.24
+Table 5: Average accuracy with different percentages of target data compared with previous SOTA LaGTran [27]. LAGUNA to obtains SOTA accuracy with $20\%$ (in green).
+Alignment in Absolute and Realtive Coordinates.
+Figure 4 shows t-SNE plots and related Maximum Mean Discrepancy (MMD) scores for $1000$ randomly selected samples from the GeoImnet validation set, where source is USA and target is Asia.
+Notably, features are separated in absolute coordinates, while they blend together (low MMD value) in relative coordinates thanks to the structure-preserving loss $\mathcal{L}_{S}$ and domain-specific learnable anchors.
+This highlights that semantic alignment can be achieved without forcing representations into a shared space, allowing each domain to preserve its unique characteristics while adhering to a common structure.
+Figure 4: t-SNE plots for 1000 randomly selected Source and Target samples from GeoImnet, with respective MMD [28] scores. Relative coordinates (right) are better aligned than absolute (left).
+Figure 5: Ablation with BLIP-1 and BLIP-2 generated captions.
+[…]
+Computational Overhead.
+While LAGUNA involves three stages and different components, its training is quite efficient: Stage 1 needs no training, and embedding 600 class labels takes 1-2 minutes ($\sim$4GB GPU memory).
+On GeoImNet, with a V100, Stage 2 requires only 9 minutes per epoch, while Stage 3 adds $<$2 minutes per epoch, which is minor considering the baseline’s 65 minutes per epoch.
+At test time, LAGUNA only requires the visual encoder, the learned anchors, the light cross-attention layer and the final MLP classifier, bringing no noticeable overhead.
 5 Conclusions
-This work introduces LAGUNA, a novel domain adaptation approach leveraging geometrical structures of semantically equivariant spaces to guide adaptation.
+We introduced LAGUNA, a novel domain adaptation approach leveraging geometrical structures of semantically equivariant spaces to guide adaptation.
 LAGUNA defines a reference representation space structure based on domain-agnostic class semantic similarities encoded by a language model, ensuring the organization of sample projections reflects this structure while preserving domain-specific characteristics.
@@
 Extensive experiments demonstrate LAGUNA’s superior performance compared to existing methods.
-This research highlights the importance of structural alignment and language-guided learning in domain adaptation.
-Future work could explore extending LAGUNA to more complex scenarios and multi-modal data.
 Supplementary Material

### v4

@@
-LAGUNA: LAnguage Guided UNsupervised Adaptation with structured spaces
+Semantically Guided Unsupervised Domain Adaptation
 Abstract
-Unsupervised domain adaptation remains a critical challenge in enabling the knowledge transfer of models across unseen domains.
+Unsupervised domain adaptation remains a critical challenge in enabling knowledge transfer of models across domains.
 Existing methods struggle to balance the need for domain-invariant representations with preserving domain-specific features.
-This is often due to the fact that alignment approaches impose the projection of samples with similar semantics close to each other in the latent space, despite drastic domain differences.
-We introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach that shifts the focus from aligning representations in absolute coordinates to aligning the relative positioning of equivalent concepts in latent spaces.
-LAGUNA defines a domain-agnostic structure on the semantic/geometric relationships between class labels in the language space and guides the adaptation, ensuring that the organization of samples in visual space reflects reference inter-class relationships while preserving domain-specific characteristics.
-Remarkably, LAGUNA surpasses previous work in $18$ different adaptation scenarios across four diverse image and video datasets with average accuracy improvements of up to $+3.32\%$ on DomainNet, $+5.75\%$ on GeoPlaces, $+4.77\%$ on GeoImnet, and $+1.94\%$ mean class accuracy improvement on EgoExo4D.
+This is often caused by alignment approaches that impose the projection of different domain samples close to each other in latent spaces, despite drastic differences.
+We introduce LAGUNA - LAnguage Guided UNsupervised Adaptation, a novel approach that shifts the focus from aligning representations in absolute coordinates to aligning the relative positioning of equivalent concepts in latent spaces.
+LAGUNA defines a domain-agnostic structure on the geometric relationships between class labels in the language space and guides the adaptation, ensuring that the organization of samples in visual space reflects reference inter-class relationships while preserving domain-specific characteristics.
+Remarkably, LAGUNA surpasses previous work in $18$ different adaptation scenarios across four diverse image and video datasets with average accuracy improvements of up to $3.32\%$ on DomainNet, $5.75\%$ on GeoPlaces, $4.77\%$ on GeoImnet, and $1.94\%$ mean class accuracy improvement on EgoExo4D.
 1 Introduction
-Domain shift challenges trained models to generalize across scenarios with differing distributions and presents a significant hurdle for supervised learning in computer vision [44].
+Domain shift challenges trained models to generalize across scenarios with differing distributions and presents a significant hurdle for supervised learning in computer vision (Luo et al., 2018).
 While fine-tuning with labeled data from the target domain can mitigate this problem, obtaining such labels often proves prohibitive.
-Unsupervised domain adaptation (UDA) [4] offers a compelling alternative, enabling knowledge transfer to novel domains without relying on expensive labeled data, which has garnered significant attention [72], promising cost-effective solutions for real-world applications prone to domain shift.
-Figure 1: Top: Existing Unsupervised Domain Adaptation approaches align source and target spaces in absolute coordinates, which may overlook the peculiarities of the distinct domains resulting in partial alignment.
-Bottom: LAGUNA aligns source (left) and target (right) spaces in relative terms, allowing data points to be distinct in absolute coordinates (e.g., circles in source and target), while $\theta_{i}^{s}$, $\theta_{i}^{t}$, their angles with other data points in the original spaces, match $\theta_{i}^{l}$, those of the reference domain-agnostic structure derived from language ($\theta_{i}^{s}$ $\sim\theta_{i}^{l}\sim$ $\theta_{i}^{t}$), encouraging the development of similar geometric-semantic relations.
+Unsupervised domain adaptation (UDA) (Bousmalis et al., 2016) offers a compelling alternative, enabling knowledge transfer to novel domains without relying on expensive labeled data, which has gained significant attention (Wilson and Cook, 2020), promising cost-effective solutions for real-world applications prone to domain shift.
 The typical UDA setting considers the availability of a labeled source domain and an unlabeled target domain.
 In general, source and target domains are semantically equivalent but drawn from distinct data distributions, e.g., real images versus clipart of chairs, TVs, and mugs.
-Thus, the main challenge of UDA is to effectively mitigate the distribution shift between domains, which is often addressed by reducing the distribution discrepancy of source and target representation spaces either by minimizing some discrepancy measure [21, 60, 65], using adversarial learning [70, 69, 34, 79, 51, 32, 42], aligning data around centroids [34], or leveraging pseudo-labels [27, 34].
+Thus, the main challenge of UDA is to effectively mitigate the distribution shift between domains, which is often addressed by reducing the distribution discrepancy of source and target representation spaces either by minimizing some discrepancy measure (Saito et al., 2018b; Tan et al., 2020), using adversarial learning (Wei et al., 2021; Li et al., 2020), aligning data around centroids (Li et al., 2021), or leveraging pseudo-labels (Kalluri et al., 2024; Li et al., 2021).
 These methods aim to align source and target representation spaces in a shared coordinate system, pushing feature vectors of equivalent semantic concepts close to each other in the embedding space, which may happen at the expense of the representation power of individual domain-specific spaces.
 For instance, bright colors and rounded shapes might be important to encode for representing a clipart, while nuanced shadows, reflections, and textures might be important to represent a real image.
-As a result, the aligned space may become overly generic, correctly encoding only a subset of the data (see Figure 1(top)).
-Recent work [48] showed that semantically equivariant representation spaces of similarly trained neural networks exhibit distinct representation spaces with matching geometrical structures.
-For instance, two data points $(x_{1},x_{2})$ may be mapped to distinct vector pairs in the two representation spaces $(v_{1}^{1},v_{2}^{1})$ and $(v_{1}^{2},v_{2}^{2})$ (e.g., $||v_{1}^{1}-v_{1}^{2}||_{2}>>0$ and $||v_{2}^{1}-v_{2}^{2}||_{2}>>0$), while angles between the two pairs will be similar in their own representation spaces (e.g., $\angle(v_{1}^{1},v_{2}^{1})\sim\angle(v_{1}^{2},v_{2}^{2})$) [^1] The reader is referred to [48] for empirical verifications of such claim..
+As a result, the aligned space may become overly generic, correctly encoding only a subset of the data (see Figure 1(left)).
+Recent work (Moschella et al., 2023) showed that semantically equivariant representation spaces of similarly trained neural networks exhibit distinct representation spaces with matching geometrical structures.
+For instance, two data points $(x_{1},x_{2})$ may be mapped to distinct vector pairs in the two representation spaces $(v_{1}^{1},v_{2}^{1})$ and $(v_{1}^{2},v_{2}^{2})$ (e.g., $v_{1}^{1}-v_{1}^{2}||_{2}>>0$ and $v_{2}^{1}-v_{2}^{2}||_{2}>>0$), while angles between the two pairs will be similar in their own representation spaces (e.g., $\angle(v_{1}^{1},v_{2}^{1})\sim\angle(v_{1}^{2},v_{2}^{2})$) (Moschella et al., 2023).
 This suggests that pushing representation spaces to overlap in absolute coordinates, as done in current domain adaptation approaches, is not necessary to obtain equivariant representation spaces.
-Based on this observation, we introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach to domain adaptation which guides source and target spaces to develop semantic-geometric inter-relationships reflecting the structure of a reference space (Figure 1(bottom)).
-As shown in recent works, language can provide a semantic space agnostic to the nuances of visual observations enabling robust zero-shot generalization [56, 10] and supporting domain robustness [13, 68, 23, 47, 27], hence we choose language to build our reference structured space in LAGUNA.
-This assumes the availability for both source and target samples of natural language descriptions, which can be generated from available captioning models [33] or collected from the web at a fraction of the cost of typical human labeling procedures [27].
+Figure 1: Left: Existing UDA approaches align source and target spaces in absolute coordinates, potentially overlooking domain-specific characteristics and resulting in partial alignment.
+Right: LAGUNA aligns spaces in relative terms, preserving distinct absolute coordinates (e.g., circles in source and target) while matching angles $\theta_{i}^{s}$, $\theta_{i}^{t}$ between data points to a reference structure $\theta_{i}^{l}$ ($\theta_{i}^{s}$ $\sim\theta_{i}^{l}\sim$ $\theta_{i}^{t}$), encouraging similar geometric-semantic relations.
+Based on this observation, we introduce LAGUNA - LAnguage Guided UNsupervised Adaptation, a novel approach to UDA which guides source and target spaces to develop semantic-geometric inter-relationships reflecting the structure of a reference space (Fig.
+1 (right)).
+As shown in recent works, language can provide a semantic space agnostic to the nuances of visual observations, enabling robust zero-shot generalization (Radford et al., 2021) and supporting domain robustness (Dunlap et al., 2022; Wang et al., 2024), hence we choose language to build our reference structured space in LAGUNA.
+This assumes the availability of both source and target samples of text descriptions, which can be generated from captioning models (Li et al., 2023) or collected from the web at a fraction of the cost of human labeling (Kalluri et al., 2024).
 LAGUNA employs a 3-stage approach to structurally align source and target representation spaces while allowing domains to preserve typical patterns (Figure 2).
-In Stage 1, textual class labels are mapped to a domain-agnostic reference space representing their relationships at the conceptual level.
-In Stage 2, a language model is trained to map textual captions to the reference latent space, providing pseudo-labels for target samples.
+In Stage 1, textual class labels are mapped to a domain-agnostic reference space representing their semantic relationships.
+In Stage 2, a language model is trained to map captions to the reference latent space, providing pseudo-labels for target samples.
 In Stage 3, a cross-domain classifier is trained to encourage domain-specific representations to follow the reference structure.
-Experiments demonstrate the superiority of LAGUNA over existing state-of-the-art methods across $18$ different domain adaptation scenarios sourced from four diverse image and video datasets, with gains of up to $+3.32\%$ on DomanNet [52], $+5.75\%$ on GeoPlaces [26], $+4.77\%$ on GeoImnet [26] and $+1.94\%$ mean per class accuracy on EgoExo4D [20].
-We further report ablations to analyze the specific contributions of our design choices.
-In sum, our main contributions are as follows: 1) we investigate the use of relative representations for UDA, showing its advantages with respect to the current absolute alignment paradigm; 2) we propose LAGUNA, a three-stage method which learns a cross-domain classifier where individual source and target spaces are distinct yet aligned to a reference structure derived from language; 3) through extensive ablations and comparisons with state-of-the-art models, we show the superiority of the proposed approach.
+Experiments demonstrate LAGUNA’s superiority over existing SOTA methods across $18$ different domain adaptation scenarios sourced from four diverse image and video datasets, with gains of up to $3.32\%$ on DomanNet (Peng et al., 2018), $5.75\%$ on GeoPlaces (Kalluri et al., 2023), $4.77\%$ on GeoImnet (Kalluri et al., 2023) and $1.94\%$ mean per class accuracy on EgoExo4D (Grauman et al., 2023).
+We further report ablations to analyze the specific contributions of our design choices and compare LAGUNA’s performance against SOTA zero-shot massive MLLMs.
+In sum, our main contributions are: 1) we investigate using relative representations for UDA, showing its advantages w.r.t absolute alignment; 2) we propose LAGUNA, a method which learns a cross-domain classifier where source and target spaces are distinct yet aligned to a reference structure; 3) through extensive ablations and comparisons with SOTA, we show the superiority of LAGUNA.
 2 Related Work
-2.1 UDA in Computer Vision
-UDA seeks to transfer knowledge from a labeled source domain to an unlabeled target domain [2, 15, 42].
-This is tackled through various approaches, prominently, discrepancy-based methods that minimize distribution differences using techniques like MMD [65, 41, 63, 28, 60, 32], and adversarial learning [4, 66, 59, 6, 70, 69, 42, 15, 66, 40].
-Other approaches also investigated class-conditional distributions alignment [51, 45, 74, 29, 70], clustering similar instances across domains [9, 50, 38, 24], instance-specific adaptation [62, 25, 67], and self-training leveraging pseudo-labeling to improve target domain performance [14, 38, 64, 54, 27].
-More recently, transformer-based approaches utilize patch composition for domain-invariant representations through intermediate domains [79].
-Video domain adaptation tackles the unique challenges of temporal dynamics and consistency in videos [8, 71, 58] with a particular focus on adapting models across the exocentric and egocentric domains [55, 22, 35, 76].
-Existing approaches sought to align source and target representation spaces in absolute coordinates, often falling short in bridging complicated forms of domain shift [26, 54].
+UDA in Computer Vision.
+UDA seeks to transfer knowledge from a labeled source domain to an unlabeled target domain (Ben-David et al., 2006; Ganin and Lempitsky, 2015; Long et al., 2018).
+This is tackled through various approaches, prominently, discrepancy-based methods that minimize distribution differences using techniques like MMD (Tan et al., 2020; Long et al., 2017b; Sun and Saenko, 2016; Li et al., 2020), and adversarial learning (Bousmalis et al., 2016; Saito et al., 2018a; Chen et al., 2019a; Wei et al., 2021; Long et al., 2018; 2017a).
+Other approaches also investigated class-conditional distributions alignment (Luo et al., 2019; Xie et al., 2018; Wei et al., 2021), clustering similar instances across domains (Deng et al., 2019; Liu et al., 2021a; Kalluri and Chandraker, 2022), instance-specific adaptation (Sharma et al., 2021; Kalluri et al., 2022; Wang et al., 2022), and self-training leveraging pseudo-labeling to improve target domain performance (French et al., 2018; Liu et al., 2021a; Sun et al., 2022; Kalluri et al., 2024).
+More recently, transformer-based approaches utilize patch composition for domain-invariant representations through intermediate domains (Zhu et al., 2023).
+Video domain adaptation tackles the unique challenges of temporal dynamics and consistency in videos (Chen et al., 2019b; Wei et al., 2023; Sahoo et al., 2021) with a particular focus on adapting models across the exocentric and egocentric domains (Quattrocchi et al., 2023; Huang et al., 2024; Xue and Grauman, 2023).
+Existing approaches sought to align source and target representation spaces in absolute coordinates, often falling short in bridging complicated forms of domain shift (Kalluri et al., 2023; Prabhu et al., 2022).
 Contrarily, LAGUNA aligns representations in relative terms, encouraging source and target visual spaces to share a similar structure as a reference space derived from language while allowing them to encode domain-specific peculiarities.
-Figure 2: The 3 stages of LAGUNA.
-In Stage 1, a domain-agnostic semantic structure in the form of a set of anchors $\mathcal{A}$ is defined.
-In Stage 2, a language model $\mathcal{G}$ is trained on source captions-label pairs to provide pseudo-labels for the target domain.
-The model is trained to be structurally aligned to anchors ${}^{*}\mathcal{A}$ with a structural loss $\mathcal{L}_{S}$ and to predict class labels with a cross-entropy loss $\mathcal{L}_{CE}$.
-In Stage 3, a cross-domain visual classifier is trained on labeled source and unlabeled target images.
-A visual encoder $\mathcal{V}$ is used to represent target and source samples as $g_{i}^{t}$ and $g_{i}^{s}$, respectively.
-The cosine similarities $\mathbf{r}^{g_{i}^{t}}$ between $g_{i}^{t}$ and learnable target anchors $\mathcal{A}_{t}$ are bound to be alike to the similarities $\mathbf{r}^{z_{i}^{t}}$ between textual representation of the target caption $z_{i}^{t}$ predicted via $\mathcal{G}$ and ${}^{*}\mathcal{A}$ by means of $\mathcal{L}_{S}$.
-A similar processing is applied to $g_{i}^{s}$, the textual representation of the ground truth class ${}^{*}A[y_{i}^{S}]$ and $\mathcal{A}_{s}$.
-A regularization loss $\mathcal{L}_{Reg}$ is used to avoid collapse during anchor learning.
-Representations $g_{i}^{t}$ and $g_{i}^{s}$ are hence grounded into the learnable $\mathcal{A}_{s}$ through a novel Cross-Domain Attention layer taking visual representations as queries, source/target anchors as keys and source anchors as values.
-Classification is finally achieved with an MLP classifier supervised via $\mathcal{L}_{CE}$ using ground truth labels in the source domain and pseudo-labels in the target domain.
-2.2 Language Guidance in Vision UDA
-Vision-language models like CLIP [56] have shown promising zero-shot transfer capabilities [10].
-However, when fine-tuned to a specific scenario [53, 1], they often lose the ability to generalize to new domains [30, 73].
-To address this issue, recent works leveraged textual information to bridge the gap between domains [13, 68, 23, 47, 18, 19].
-In particular, the seminal work of [27] proposed to use rich textual captions to provide pseudo-labels for the target domain and addressed UDA by training a joint classifier on source and target domains.
-Similarly to [27], we generate pseudo-labels from textual captions, but, rather than seeking to align the visual representation spaces of source and target domains to language in an absolute reference frame, LAGUNA uses language to guide a relative alignment between source and target visual spaces.
-2.3 Relative Encodings
-Recent work [48] has shown that equivalent latent spaces of similarly trained networks tend to be misaligned in absolute terms but share similar internal geometric relationships.
-As a result, representing data points with relative encodings, obtained as vectors of cosine similarity values of a data point with a predefined set of anchors, allows to perform zero-shot model stitching [48], translate representation spaces across models [46] or modalities [49].
-As shown in [5], predefined invariances can be incorporated into the learned representation to enable specific forms of relative representations.
-In [11] relative encodings were used to tackle the downstream task of action anticipation.
+Language Guidance in Vision UDA.
+Vision-language models like CLIP (Radford et al., 2021) have shown promising zero-shot transfer capabilities (Devillers et al., 2021).
+However, when fine-tuned to a specific scenario (Pham et al., 2021; Andreassen et al., 2021), they lose the ability to generalize to new domains (Kumar et al., 2022; Wortsman et al., 2021).
+To address this, recent works leveraged textual information to bridge the gap between domains (Dunlap et al., 2022; Wang et al., 2024; Huang et al., 2023; Min et al., 2022; Gokhale et al., 2020; Goyal et al., 2022).
+In particular, the seminal work of (Kalluri et al., 2024) proposed to use textual captions to provide pseudo-labels for the target domain and addressed UDA by training a joint classifier on source and target domains.
+Similarly to Kalluri et al.
+(2024), we generate pseudo-labels from textual captions, but, rather than seeking to align the visual representation spaces of source and target domains to language in an absolute reference frame, LAGUNA uses language to guide a relative alignment between source and target visual spaces.
+Relative Encodings.
+Recent work (Moschella et al., 2023) has shown that equivalent latent spaces of similarly trained networks tend to be misaligned in absolute terms but share similar internal geometric relationships.
+As a result, using relative encodings, obtained as similarity values of data points w.r.t a predefined set of anchors, allows to perform zero-shot model stitching (Moschella et al., 2023), translate representation spaces across models (Maiorca et al., 2023) or modalities (Norelli et al., 2023).
+As shown in (Cannistraci et al., 2024), predefined invariances can be incorporated into the learned representation to enable specific forms of relative representations.
+In (Diko et al., 2024) relative encodings were used to tackle the downstream task of action anticipation.
 LAGUNA builds on relative encodings to represent the semantic inter-relations between classes in a domain-agnostic language-based reference space and support the development of aligned yet specialized source and target domains.
+Figure 2: The 3-stage architecture.
+(1) We define domain-agnostic semantic anchors $\mathcal{A}$.
+(2) A language model $\mathcal{G}$ generates pseudo-labels for target data, trained with structural loss $\mathcal{L}_{S}$ and cross-entropy loss $\mathcal{L}_{CE}$.
+(3) A visual classifier is trained using an encoder $\mathcal{V}$ to extract features $g_{i}^{s}$ and $g_{i}^{t}$.
+We align visual-anchor similarities with text-anchor similarities using $\mathcal{L}_{S}$, learnable anchors ($\mathcal{A}_{s}$, $\mathcal{A}_{t}$), and textual representations ($z_{i}^{t}$, $^{*}A[y_{i}^{S}]$).
+A Cross-Domain Attention layer grounds visual features using $\mathcal{A}_{s}$, and an MLP classifier is trained with $\mathcal{L}_{CE}$ and regularized by $\mathcal{L}_{Reg}$.
 3 Method
-LAGUNA comprises a 3-stage process, as illustrated in Fig.
+LAGUNA works in 3 stages, as shown in Fig.
 2.
-In Stage 1, a language-based reference structure space is defined.
-We represent this structure as a set of anchors, which we will use as a reference to guide representation learning.
-In Stage 2, a text-supervisor language model is trained to map image/video captions to class categories, providing pseudo-labels for the unlabeled target domain.
-The language model is also trained to keep internal text embeddings aligned to the domain-agnostic structure defined in Stage 1.
-In Stage 3, a cross-domain visual classifier is trained to categorize source and target samples by learning domain-specific action anchors that are structurally similar to the domain-agnostic structure defined in Stage 1.
+First, we create a language-based reference structure using anchor points to guide representation learning.
+Second, we train a language model to map image/video captions to class categories, generating pseudo-labels for unlabeled target data while keeping text embeddings aligned to the reference structure.
+Finally, we train a cross-domain visual classifier that learns domain-specific action anchors structured similarly to our reference framework.
 3.1 Problem Setup
-We follow the formulation of [27] and assume a labeled source domain $\mathcal{X}_{s}:\{(x_{i}^{s},y_{i}^{s},l_{i}^{s})\}_{i=1}^{N_{s}}$, where samples $x_{i}^{s}$ are paired both with labels $y_{i}^{s}$ and language descriptions $l_{i}^{s}$, whereas the target domain $\mathcal{X}_{t}:\{x_{i}^{t},l_{i}^{t}\}_{i=1}^{N_{t}}$ contains unlabeled samples $x_{i}^{t}$ paired with language descriptions $l_{i}^{t}$.
-These textual descriptions can be readily obtained from associated metadata or generated using image-to-text models [33].
+We follow the formulation of Kalluri et al.
+(2024) and assume a labeled source domain $\mathcal{X}_{s}:\{(x_{i}^{s},y_{i}^{s},l_{i}^{s})\}_{i=1}^{N_{s}}$, where samples $x_{i}^{s}$ are paired with labels $y_{i}^{s}$ and captions $l_{i}^{s}$, whereas the target domain $\mathcal{X}_{t}:\{x_{i}^{t},l_{i}^{t}\}_{i=1}^{N_{t}}$ contains unlabeled samples $x_{i}^{t}$ paired with language descriptions $l_{i}^{t}$.
+These textual descriptions can be readily obtained from associated metadata or generated using image-to-text models (Li et al., 2023).
 $N_{s}$ and $N_{t}$ represent the number of source and target samples, respectively, and the two domains share the same high-level semantics and number of classes $N_{c}$.
 3.2 Stage 1 - Reference Structure Definition
-Assuming shared classes across domains, in Stage 1 (Fig.
-2), we construct a language-based, domain-agnostic reference structure in the form of a set of vectors $\mathcal{A}\in\mathbb{R}^{N_{c}\times D_{l}}$.
-Each vector is obtained by encoding the $N_{c}$ class label names through a pre-trained SentenceTransformer model with output dimensionality $D_{l}$ trained for semantic similarity [57].
-Following literature on relative representations [48], we treat the vectors in $\mathcal{A}$ as a reference set of anchors.
-Given a vector $v\in\mathbb{R}^{D_{l}}$, we define its relative encoding with respect to anchors $\mathcal{A}$ as $\mathbf{r}^{v}=rel(v,\mathcal{A})=[\cos(v,\mathcal{A}[1]),\ldots,\cos(v,\mathcal{A}[N_{c}])]$, where $\cos(\cdot,\cdot)$ is the cosine similarity (i.e., the cosine of the angle $\theta$ between two vectors), and $\mathcal{A}[i]$ is the anchor in $\mathcal{A}$ associated to class $i$.
-Intuitively, vector $\mathbf{r}^{v}$ encodes the geometrical/semantic relationships between vector $v$ and the reference language anchors $\mathcal{A}$.
-Following this logic, the set of all reference affinities is defined as $\mathbf{r}^{\mathcal{A}}=[rel(\mathcal{A}[i],\mathcal{A}),\dots,rel(\mathcal{A}[N_{c}],\mathcal{A})]$ where each class anchors $\mathcal{A}[i]$ is represented according to its relationship with respect to the other anchors in $\mathcal{A}$.
-During the training of the cross-domain visual classifier (Stage 3), these encodings are used to enforce the learned latent space to follow a similar structure as the one induced by $\mathcal{A}$.
+Assuming shared classes across domains, Stage 1 (Fig.
+2) builds a language-based, domain-agnostic reference structure as a set of vectors $\mathcal{A}\in\mathbb{R}^{N_{c}\times D_{l}}$.
+We obtain each vector by encoding the $N_{c}$ class label names using a pre-trained SentenceTransformer model with output dimensionality $D_{l}$ trained for semantic similarity (Reimers and Gurevych, 2019).
+Following the relative representations literature (Moschella et al., 2023), we treat these vectors in $\mathcal{A}$ as reference anchors.
+For any vector $v\in\mathbb{R}^{D_{l}}$, we define its relative encoding with respect to anchors $\mathcal{A}$ as $\mathbf{r}^{v}=rel(v,\mathcal{A})=[\cos(v,\mathcal{A}[1]),\ldots,\cos(v,\mathcal{A}[N_{c}])]$, where $\cos(\cdot,\cdot)$ represents cosine similarity and $\mathcal{A}[i]$ is the anchor for class $i$.
+This vector $\mathbf{r}^{v}$ captures the geometric and semantic relationships between $v$ and our reference language anchors $\mathcal{A}$.
+We define the complete set of reference affinities as $\mathbf{r}^{\mathcal{A}}=[rel(\mathcal{A}[1],\mathcal{A}),\dots,rel(\mathcal{A}[N_{c}],\mathcal{A})]$, where each class anchor $\mathcal{A}[i]$ is represented by its relationships to all other anchors in $\mathcal{A}$.
+During Stage 3’s cross-domain visual classifier training, these encodings enforce the learned latent space to follow the same structure induced by $\mathcal{A}$.
 3.3 Stage 2: Training of the Language Supervisor
-Similar to [27], we train a language model to map image captions to semantic labels, with the aim to provide pseudo-labels for target samples, which are associated with captions but not class labels.
-Differently from [27], we use the language supervisor both to provide pseudo-labels for target samples, useful to train the classifier, and textual representations semantically structured as the domain-agnostic anchors $\mathcal{A}$, useful to encourage alignment to the reference structure.
-Hence, rather than training a regular classifier, we directly supervise a language model $\mathcal{G}$ to provide representations that are 1) geometrically aligned to $\mathcal{A}$ and 2) suitable for predicting class labels.
-Specifically, given a pair of source caption and corresponding label $(l_{i}^{s},y_{i}^{s})$, $\mathcal{G}$ processes $l_{i}^{s}$ to produce a vector representation $z^{s}_{i}$:
-$z_{i}^{s}=\mathcal{G}(l_{i}^{s}).$ (1)
+Similar to (Kalluri et al., 2024), we train a language model ($\mathcal{G}$) from captions to provide pseudo-labels for target samples, which have no class labels.
+In addition, we also use $\mathcal{G}$ to learn textual representations semantically structured as the domain-agnostic anchors $\mathcal{A}$, useful to encourage alignment to the reference structure.
+Hence, rather than training a regular classifier, we directly supervise $\mathcal{G}$ to provide representations that are 1) geometrically aligned to $\mathcal{A}$ and 2) suitable for predicting class labels.
+Specifically, given a pair of source caption-label $(l_{i}^{s},y_{i}^{s})$, $\mathcal{G}$ processes $l_{i}^{s}$ to produce a vector representation $z^{s}_{i}=\mathcal{G}(l_{i}^{s})$.
 Next, we encourage the vector representation $z_{i}^{s}$ to be geometrically aligned to $\mathcal{A}[y_{i}^{s}]$ corresponding to ground truth action $y_{i}^{s}$.
@@
 $\mathcal{L}_{S}=|\mathbf{r}^{z_{i}^{s}}-\mathbf{r}^{y_{i}^{s}}|.$ (1)
-This loss encourages the encodings $z_{i}^{s}$ of text descriptions $l_{i}^{s}$ associated with labels $y_{i}^{s}$ to preserve the same geometrical associations as their corresponding anchor, encouraging the latent space learned by $\mathcal{G}$ to mirror the structure defined by $\mathcal{A}$.
+This loss encourages the encodings $z_{i}^{s}$ of text descriptions $l_{i}^{s}$ associated with $y_{i}^{s}$ to preserve the same geometrical associations as $\mathcal{A}[y_{i}^{s}]$, encouraging the latent space learned by $\mathcal{G}$ to mirror the structure defined by $\mathcal{A}$.
 To further favor alignment to $\mathcal{A}$, rather than employing a classification head, we predict class probabilities directly by Softmax-normalizing relative encodings:
@@
 3.4 Stage 3: Cross-Domain Visual Classifier
-Stage 3 trains a cross-domain visual classifier aligning representations extracted through a visual encoder $\mathcal{V}$ to the structure imposed by the domain-agnostic language anchors $\mathcal{A}$.
+Stage 3 trains a cross-domain visual classifier aligning representations extracted through a visual encoder $\mathcal{V}$ to the structure imposed by $\mathcal{A}$.
 We allow each domain to develop its own latent space, but we encourage both spaces to be aligned to two sets of learnable anchors $\mathcal{A}_{t}\in\mathbb{R}^{N_{c}\times D_{v}}$ and $\mathcal{A}_{s}\in\mathbb{R}^{N_{c}\times D_{v}}$ specific to the target and source domain respectively.
-We further ensure that the structures of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are aligned to the structure of $\mathcal{A}$ through the supervision provided by ground truth labels $y_{i}^{s}$ in the source domain and the text supervision provided by $\mathcal{G}$ in the target domain in the form of the predicted vector $z_{i}^{t}$ and the predicted pseudo-label $\overline{y}_{i}^{t}$.
+We further ensure that the structures of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are aligned to that of $\mathcal{A}$ through supervision provided by $y_{i}^{s}$ in the source domain and the text supervision provided by $\mathcal{G}$ in the target domain in the form of $z_{i}^{t}$ and $\overline{y}_{i}^{t}$.
 The source and target datasets are merged ($\hat{\mathcal{X}}=\mathcal{X}_{t}+\mathcal{X}_{s}$) and used for training with a total of $M=N_{s}+N_{t}$ samples.
-Source samples comprise an image, caption, and label triplet $(x_{i}^{s},l_{i}^{s},y_{i}^{s})$, whereas target samples $(x_{i}^{t},l_{i}^{t},\overline{y}_{i}^{t})$ replace label $y_{i}^{s}$ with the pseudo-label $\overline{y}_{i}^{t}$.
+Source samples comprise $(x_{i}^{s},l_{i}^{s},y_{i}^{s})$ triplet, whereas target samples $(x_{i}^{t},l_{i}^{t},\overline{y}_{i}^{t})$ use $\overline{y}_{i}^{t}$ in absence of $y_{i}^{t}$.
 Target and source images are processed by $\mathcal{V}$ to obtain latent representations $g_{i}^{t}$ and $g_{i}^{s}$, while target captions $l_{i}^{t}$ are processed by $\mathcal{G}$ to obtain latent representations $z_{i}^{t}$ as follows:
 $g_{i}^{t}=\mathcal{V}(x_{i}^{t}),\hskip 14.22636ptz_{i}^{t}=\mathcal{G}(l_{i}^{t}),\hskip 14.22636ptg_{i}^{s}=\mathcal{V}(x_{i}^{s}).$ (4)
-These representations and the anchors are used to achieve two main training objectives: 1) structure learning and 2) classification training.
 3.4.1 Structure Learning
 To ensure the learned visual representations adhere to the structure defined by $\mathcal{A}$, LAGUNA employs a supervised learning approach based on relative encodings.
-For target domain samples (without ground truth labels), we first compute the relative encoding $\mathbf{r}^{z_{i}^{t}}=rel(z_{i}^{t},\mathcal{A})$.
-For source domain samples, we compute the relative encoding of the anchor associated to the ground truth class $\mathcal{A}[y_{i}^{s}]$: $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$.
+For target samples, we first compute $\mathbf{r}^{z_{i}^{t}}=rel(z_{i}^{t},\mathcal{A})$.
+For source samples, we compute the relative encoding of the anchor associated with the ground truth class $\mathcal{A}[y_{i}^{s}]$: $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$.
 These encodings represent the relative positions of textual representations with respect to the reference space $\mathcal{A}$ and are not trainable as $\mathcal{G}$ is fixed.
@@
 $\mathcal{L}_{S}=\begin{cases}|\mathbf{r}^{g_{i}^{t}}-\mathbf{r}^{z_{i}^{t}}|,&\textit{if Target Domain}\\ |\mathbf{r}^{g_{i}^{s}}-\mathbf{r}^{y_{i}^{s}}|,&\textit{otherwise}\end{cases}.$ (5)
-This loss encourages visual representations to maintain relationships with their respective domain-specific visual anchors that mirror the relationships expressed by the domain-invariant anchors $\mathcal{A}$.
+This loss encourages $g_{i}^{t}$ and $g_{i}^{s}$ to maintain relationships with their respective domain-specific visual anchors that mirror those expressed by $\mathcal{A}$.
 This process supervises both the learnable anchors $\mathcal{A}_{s},\mathcal{A}_{t}$ and the visual encoder $\mathcal{V}$.
-However, since this loss alone might lead to the collapse of anchor representations into a small region, hindering effective classification, we introduce a volume (or spread) regularization loss.
-This loss treats each set of anchors as a multidimensional parallelotope whose volume in the latent space can be measured by the determinant of its Gram matrix.
+However, since this loss alone might lead to the collapse of anchor representations into a small region, hindering classification decision boundaries, we introduce a volume (spread) regularization loss.
+This loss treats each set of anchors as a multidimensional parallelotope whose volume in the latent space is measured by the determinant of its Gram matrix.
 Given the Gram matrices for the three sets of anchors:
 $\gamma=\mathcal{A}\mathcal{A}^{T},\hskip 14.22636pt\gamma_{s}=\mathcal{A}_{s}\mathcal{A}_{s}^{T},\hskip 14.22636pt\gamma_{t}=\mathcal{A}_{t}\mathcal{A}_{t}^{T},$ (6)
-we encourage the volume of each domain-specific parallelotope to be approximately equal to the volume of the reference anchors $\mathcal{A}$ with the following regularization loss:
-$\mathcal{L}_{Reg}$ $=$ $|logDet(\gamma_{t})-logDet(\gamma)|+$ (8)
-$|logDet(\gamma_{s})-logDet(\gamma)|,$
-where we use log-determinants for numerical stability.
-This regularization loss ensures taht visual anchors occupy a volume similar to the one of $\mathcal{A}$, preventing collapse and promoting better class separation in the representation space.
+we encourage the volume of each domain-specific parallelotope to be approximately equal to that of $\mathcal{A}$ with the following loss:
+$\mathcal{L}_{Reg}=|logDet(\gamma_{t})-logDet(\gamma)|+|logDet(\gamma_{s})-logDet(\gamma)|.$ (7)
+We use log-determinants for numerical stability. This regularization loss ensures that visual anchors occupy a volume similar to that of $\mathcal{A}$, preventing collapse in the representation space.
 3.4.2 Classification Training
 We perform classification training concurrently with structure learning.
-To ensure that source and target visual representations are grounded in the structure imposed by $\mathcal{A}$, while still being free to capture domain-specific nuances, we propose a novel cross-domain attention layer that aims to ground domain-specific visual representations $g_{i}^{t}$ and $g_{i}^{s}$ to the structure of source anchors $\mathcal{A}_{s}$.
+To ensure that source and target visual representations are grounded in the structure imposed by $\mathcal{A}$, while still being free to capture domain-specific nuances, we propose a novel cross-domain attention layer that aims to ground $g_{i}^{t}$ and $g_{i}^{s}$ to the structure of source anchors $\mathcal{A}_{s}$.
 We include two Cross-Domain attention layers sharing the same weights for the target and source branches of LAGUNA (See Fig.
@@
 Finally, an MLP classification head processes $f_{i}^{t}$ and $f_{i}^{s}$ to output predictions $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{s}$.
-The MLP has shared weights across domains and is optimized with a cross-entropy loss using the ground truth label $y_{i}^{s}$ for source examples and the pseudo-label $\overline{y}_{i}^{t}$ predicted by $\mathcal{G}$ in the case of target samples.
+The MLP has shared weights across domains and is optimized with a cross-entropy loss using $y_{i}^{s}$ for source examples and $\overline{y}_{i}^{t}$ in the case of target samples.
 The overall training objective is:
-$\small\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(\hat{y_{i}},y_{i}^{*})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{*},\mathbf{r}^{g_{i}})+\lambda_{3}\mathcal{L}_{Reg}(\gamma,\gamma_{*}),$ (11)
+$\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(\hat{y_{i}},y_{i}^{*})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{*},\mathbf{r}^{g_{i}})+\lambda_{3}\mathcal{L}_{Reg}(\gamma,\gamma_{*}),$ (10)
 where $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ are hyperparameter weights to calibrate the magnitude of each loss value, and ’*’ signifies that the argument is domain-dependent [^1] For example, for $\mathcal{L}_{CE}$ the ground truth can either be $y_{i}^{s}$ or $\overline{y}_{i}^{t}$..
 4 Experiments
+4.1 Experimental Setup
+Datasets.
+We evaluate LAGUNA on four comprehensive UDA benchmarks: DomainNet (Peng et al., 2018), GeoPlaces (Kalluri et al., 2023) GeoImnet (Kalluri et al., 2023), and Ego2Exo (Kalluri et al., 2024).
+DomainNet consists of $400K$ images across $345$ classes and is used to evaluate performance in $12$ adaptation settings.
+GeoImnet and GeoPlaces, subsets of the GeoNet dataset, contain over 750K images, and focus on geographic disparities for objects (GeoImNet with $600$ classes) and scenes (GeoPlaces with $205$ classes).
+Finally, Ego2Exo is a subset of EgoExo-4D (Grauman et al., 2023) curated for domain adaptation.
+It involves a total of $9086$ videos and $24$ categories.
+DomainNet and GeoNet include per-image captions derived from BLIP-2 (Li et al., 2023) and web metadata, respectively, while Ego2Exo provides video descriptions from the original EgoExo-4D.
+Model and training details.
+We use a pre-trained SentenceTransformer (Reimers and Gurevych, 2019) for Stage 1 and a BERT base model (Sanh et al., 2019) as text supervisor (Stage 2).
+The latter is trained on the source domain of each scenario for 5 epochs with a batch size of 64, a lr of 1e-4, and the AdamW (Loshchilov, 2017) optimizer.
+The visual encoders (Stage 3) are tailored to each dataset for fair comparisons with existing approaches.
+For DomainNet, we adapt the Swin-B and ViT-B backbones (Liu et al., 2021b; Dosovitskiy et al., 2020).
+For GeoImnet and GeoPLaces, we utilize the ViT-B backbone.
+For Ego2Exo, we employ pre-extracted Omnivore-L features (Grauman et al., 2023).
+All vision models undergo joint training with the other network components for $10$ epochs with a batch size of $32$, an initial learning rate of 1e41e-4, and cosine scheduling.
+We set the loss weights to $\lambda_{1}$=1.0, $\lambda_{2}$=0.1, and $\lambda_{3}$=0.001 to normalize the magnitudes of loss components.
 […]
@@
 Source Only 63.02 49.47 60.48 70.52 56.09 52.53 70.42 65.91 54.47 73.34 60.09 48.25 60.38
-MDD [77] 52.80 41.20 47.80 52.50 42.10 40.70 54.20 54.30 43.10 51.20 43.70 41.70 47.11
-SCDA [34] 54.00 42.50 51.90 55.00 44.10 39.30 53.20 55.60 44.70 56.20 44.10 42.00 48.55
-SSRT-B [64] 69.90 58.90 66.00 75.80 59.80 60.80 73.20 70.60 62.20 71.40 61.70 55.20 65.41
-MemSAC [25] 63.49 42.14 60.32 72.33 54.92 46.14 73.46 68.04 52.75 74.42 57.79 43.57 59.11
-CDTrans [75] 66.20 52.90 61.50 72.60 58.10 57.10 72.50 69.00 59.00 72.10 62.90 53.90 63.16
-PMTrans [79] 74.10 61.10 70.00 79.30 63.70 62.70 77.50 73.80 62.60 79.80 69.70 61.20 69.63
+MDD (Zhang et al., 2019) 52.80 41.20 47.80 52.50 42.10 40.70 54.20 54.30 43.10 51.20 43.70 41.70 47.11
+SCDA (Li et al., 2021) 54.00 42.50 51.90 55.00 44.10 39.30 53.20 55.60 44.70 56.20 44.10 42.00 48.55
+SSRT-B (Sun et al., 2022) 69.90 58.90 66.00 75.80 59.80 60.80 73.20 70.60 62.20 71.40 61.70 55.20 65.41
+MemSAC (Kalluri et al., 2022) 63.49 42.14 60.32 72.33 54.92 46.14 73.46 68.04 52.75 74.42 57.79 43.57 59.11
+CDTrans (Xu et al., 2021) 66.20 52.90 61.50 72.60 58.10 57.10 72.50 69.00 59.00 72.10 62.90 53.90 63.16
+PMTrans (Zhu et al., 2023) 74.10 61.10 70.00 79.30 63.70 62.70 77.50 73.80 62.60 79.80 69.70 61.20 69.63
 UDA with Language Guidance
-TextMatch [27] 71.36 64.30 65.32 81.25 65.65 64.85 81.09 72.65 63.94 81.08 70.84 64.17 70.64
-nGramMatch [27] 68.92 59.82 63.15 76.35 61.72 62.87 76.35 69.28 62.51 76.04 68.52 60.52 67.17
-LaGTran [27] 77.30 68.25 67.35 81.31 67.03 66.81 80.78 75.62 68.08 79.23 73.80 63.44 72.41
+TextMatch (Kalluri et al., 2024) 71.36 64.30 65.32 81.25 65.65 64.85 81.09 72.65 63.94 81.08 70.84 64.17 70.64
+nGramMatch (Kalluri et al., 2024) 68.92 59.82 63.15 76.35 61.72 62.87 76.35 69.28 62.51 76.04 68.52 60.52 67.17
+LaGTran (Kalluri et al., 2024) 77.30 68.25 67.35 81.31 67.03 66.81 80.78 75.62 68.08 79.23 73.80 63.44 72.41
 LAGUNA 80.34 70.68 71.92 83.07 69.51 70.59 83.34 79.71 70.51 83.32 77.47 68.32 75.73
 Improvement +3.04 +2.43 +1.92 +1.82 +2.48 +3.78 +2.25 +4.09 +2.43 +2.24 +3.67 +4.15 +3.32
-ViT-B CLIP* [56] 72.39 60.90 66.81 81.37 60.90 66.81 81.37 72.39 66.81 81.37 72.39 69.90 70.38
-DAPrompt* [16] 73.90 65.90 70.40 84.90 65.80 70.20 84.60 73.50 69.90 84.90 73.80 65.80 73.80
-PADCLIP [31] 76.40 67.50 72.70 84.20 68.10 71.10 83.60 76.30 71.70 83.50 75.40 67.20 74.81
+ViT-B CLIP* (Radford et al., 2021) 72.39 60.90 66.81 81.37 60.90 66.81 81.37 72.39 66.81 81.37 72.39 69.90 70.38
+DAPrompt* (Ge et al., 2025) 73.90 65.90 70.40 84.90 65.80 70.20 84.60 73.50 69.90 84.90 73.80 65.80 73.80
+PADCLIP (Lai et al., 2023) 76.40 67.50 72.70 84.20 68.10 71.10 83.60 76.30 71.70 83.50 75.40 67.20 74.81
 […]
-UniMoS* [36] 77.50 68.00 72.50 86.00 68.50 72.30 85.90 77.80 72.60 85.80 77.20 68.20 76.03
+UniMoS* (Li et al., 2024) 77.50 68.00 72.50 86.00 68.50 72.30 85.90 77.80 72.60 85.80 77.20 68.20 76.03
 LAGUNA 79.53 73.18 74.24 86.04 73.54 74.98 85.81 80.24 74.43 86.57 79.89 74.11 78.54
 Improvement +2.03 +4.07 +1.74 +0.04 +5.03 +2.68 -0.09 +2.44 +1.83 +0.77 +2.69 +5.91 +2.51
-Table 1: Results on DomainNet [52] in $4$ domains: Real, Clipart, Sketch, and Painting, for 12 adaptation scenarios using the Swin-B [39] and ViT-B [12] backbones.
-The best results are in bold, second bests are underlined.‘*‘ method uses domain-aware prompts, $\dagger$ our run.
-4.1 Experimental Setup
-Datasets.
-We evaluate LAGUNA on four comprehensive UDA benchmarks: DomainNet [52], GeoPlaces [26] GeoImnet [26], and Ego2Exo [27, 20].
-DomainNet consists of $400K$ images across $345$ classes and is used to evaluate performance in $12$ domain transfer settings [70, 27].
-GeoImnet and GeoPlaces are subsets of the larger GeoNet dataset [26], contain over 750K images and focus on geographic disparities in image classification (GeoImNet with $600$ classes) and scene recognition (GeoPlaces with $205$ classes).
-Finally, Ego2Exo is a subset of EgoExo-4D [20] curated for domain adaptation.
-It involves a total of $9086$ videos and $24$ action categories.
-DomainNet and GeoNet include per-image captions derived from BLIP-2 [33] and web metadata, respectively, while Ego2Exo provides video descriptions from the original EgoExo-4D dataset.
-Model and training details.
-We use a pre-trained SentenceTransformer [57] for structure definition (Stage 1) and a BERT base model [61] as text supervisor (Stage 2).
-The latter is trained on the source domain of each scenario for 5 epochs with a batch size of 64, a lr of 1e-4, and the AdamW [43] optimizer.
-The visual encoders (Stage 3) are tailored to each dataset for fair comparisons with existing approaches.
-For DomainNet, we adapt the Swin-B and ViT-B backbones [39, 12].
-For GeoImnet and GeoPLaces, we utilize the ViT-B backbone [12] following [70, 79].
-For Ego2Exo, we employ pre-extracted Omnivore-L features [20] as in [27].
-All vision models undergo joint training with the other network components for $10$ epochs with a batch size of $32$, an initial learning rate of $41e-$ 4, and cosine scheduling.
-We set the loss weights to $\lambda_{1}$=1.0 $\lambda_{2}$=0.1 and $\lambda_{3}$=0.001 to normalize the relative magnitude of the related loss functions.
+Table 1: DomainNet dataset results for domain adaptation.
+Includes 4 domains: Real, Clipart, Sketch, and Painting, for a total of 12 domain adaptation scenarios.
+The best results are reported in bold while the second best are underlined.
+4.2 Results
+We benchmark our model against various language-free UDA methods that have reported results on the considered datasets (Zhu et al., 2023; Wei et al., 2021; Kalluri et al., 2022; Zhang et al., 2019; Chen et al., 2022; Xu et al., 2021).
+Additionally, we also compare with previous approaches using language guidance (Kalluri et al., 2024; Radford et al., 2021; Ge et al., 2025; Lai et al., 2023; Li et al., 2024).
+Following the evaluation protocols of each benchmark, we report accuracy for DomainNet, GeoImNet, and GeoPlaces, and per-class mean accuracy for Ego2Exo.
+DomainNet.
+Table 1 presents the results on DomainNet, encompassing $12$ UDA scenarios across four distinct domains: Real (R), Clipart (C), Sketch (S), and Painting (P).
+LAGUNA is compared against previous SOTA methods utilizing the Swin-B and ViT-B backbones as originally adopted by the compared methods to ensure fair comparison.
+LAGUNA significantly outperforms previous methods in 12 scenarios, with an average performance increase of $3.32\%$ for Swin-B.
+For ViT-B, LAGUNA leads in 11 of 12 scenarios, boosting performance by an average of $2.51\%$.
+Importantly, LAGUNA achieves SOTA ViT-B performance without relying on extensive CLIP pretraining and domain-aware prompting, unlike previous models (Lai et al., 2023; Li et al., 2024; Ge et al., 2025).
+[…]
+[…]
 Model Geolmnet GeoPlaces Avg.
@@
 Source Only 52.46 51.91 44.90 36.85 46.53
-CDAN [42] 54.48 53.87 42.88 36.21 46.86
-MemSAC [25] 53.02 54.37 42.05 38.33 46.94
-ToAlign [70] 55.67 55.92 42.32 38.40 48.08
-MDD [77] 51.57 50.73 42.54 39.23 46.02
-DALN [7] 55.36 55.77 41.06 40.41 48.15
-PMTrans [79] 56.76 57.60 46.18 40.33 50.22
+CDAN (Long et al., 2018) 54.48 53.87 42.88 36.21 46.86
+MemSAC (Kalluri et al., 2022) 53.02 54.37 42.05 38.33 46.94
+ToAlign (Wei et al., 2021) 55.67 55.92 42.32 38.40 48.08
+MDD (Zhang et al., 2019) 51.57 50.73 42.54 39.23 46.02
+DALN (Chen et al., 2022) 55.36 55.77 41.06 40.41 48.15
+PMTrans (Zhu et al., 2023) 56.76 57.60 46.18 40.33 50.22
 UDA with Language Guidance
-TextMatch [27] 49.68 54.82 53.06 50.11 51.92
-nGramMatch [27] 49.53 51.02 51.70 49.87 50.93
-LaGTran [27] 63.67 64.16 56.14 57.02 60.24
+TextMatch (Kalluri et al., 2024) 49.68 54.82 53.06 50.11 51.92
+nGramMatch (Kalluri et al., 2024) 49.53 51.02 51.70 49.87 50.93
+LaGTran (Kalluri et al., 2024) 63.67 64.16 56.14 57.02 60.24
 LAGUNA 67.39 69.97 61.15 63.51 65.40
-Improvement +3.72 +5.81 +5.01 +6.49 +5.26
-Table 2: Results on GeoImnet [26] and GeoPlaces [26] with 4 adaptation scenarios. Best results are in bold, whereas second best results are underlined. All methods use the ViT-B [12] backbone.
-4.2 Results
-We benchmark our model against various language-free domain adaptation methods that have reported results on the considered datasets [79, 70, 25, 77, 7, 75].
-Besides classic UDA methods, we also compare with previous approaches using language guidance [27, 56, 16, 31, 36].
-Following the evaluation protocols of each benchmark, we report accuracy for DomainNet, GeoImNet, and GeoPlaces, and per-class mean accuracy for Ego2Exo.
-DomainNet.
-Table 1 presents the results on the DomainNet dataset, encompassing $12$ domain adaptation scenarios across four distinct domains: Real (R), Clipart (C), Sketch (S), and Painting (P).
-LAGUNA is compared against previous state-of-the-art methods utilizing the Swin-B and ViT-B backbones as originally adopted by the compared methods to ensure fair comparison.
-LAGUNA significantly outperforms previous methods in 12 scenarios, with an average performance increase of $3.32\%$ for the Swin-B backbone.
-For ViT-B, LAGUNA leads in 11 of 12 scenarios, boosting performance by an average of $2.51\%$.
-Importantly, LAGUNA achieves SOTA ViT-B performance without relying on extensive CLIP pretraining and domain-specific prompting, unlike previous models [31, 36, 16].
-[…]
-[…]
-These consistent improvements across all tested datasets, with a total of 18 scenarios, underscore the effectiveness of LAGUNA with respect to both language-free and language-guided UDA approaches.
+Improvement +3.72 +5.81 +5.01 +6.49 +5.16
+Table 2: Results on GeoImnet and GeoPlaces with 4 adaptation scenarios using ViT-B. Best results are in bold, whereas second-best are underlined.
 […]
 Source Only 8.39 15.66 12.03
-TA3N [8] 6.92 27.95 17.44
-TransVAE [71] 12.06 23.34 17.70
+TA3N (Chen et al., 2019b) 6.92 27.95 17.44
+TransVAE (Wei et al., 2023) 12.06 23.34 17.70
 Zero-shot Video Recognition
-EgoVLP [37] 5.89 19.35 12.62
-LaVILA [78] 5.86 23.16 14.51
+EgoVLP (Lin et al., 2022) 5.89 19.35 12.62
+LaVILA (Zhao et al., 2022) 5.86 23.16 14.51
 UDA with Language Guidance
-TextMatch [27] 10.36 13.57 11.97
-nGramMatch [27] 11.50 15.46 13.98
-LaGTran [27] 12.34 30.76 21.55
+TextMatch (Kalluri et al., 2024) 10.36 13.57 11.97
+nGramMatch (Kalluri et al., 2024) 11.50 15.46 13.98
+LaGTran (Kalluri et al., 2024) 12.34 30.76 21.55
 LAGUNA 13.52 33.45 23.49
@@
 Improvement +1.18 +2.69 +1.94
-Table 3: Results on the Ego2Exo benchmark from EgoExo-4D [20].
-Best results are in bold, whereas second best results are underlined.
-All methods use official pre-extracted Omnivore-L features [17, 20] except EgoVLP [37] and LaVILA [78], which use Timesformer-B [3] backbone.
+Table 3: Results on the Ego2Exo benchmark. Best results are in bold, second best are underlined. All methods use Omnivore-L features except EgoVLP and LaVILA.
 4.3 Ablation Study
-We analyze the effect of the individual components of LAGUNA to provide an in-depth assessment of model performance and justify design choices.
-We investigate (a) the impact of our methodological elements, (b) the effects of different reference structures, (c) the behavior of LAGUNA with different ratios of pseudo-labels, and (d) the qualitative illustration of the alignment of relative and absolute spaces.
-More ablations are reported in the Supplementary Material.
+In this section, we conduct comprehensive ablation studies across multiple dimensions of LAGUNA: 1) methodological component analysis, 2) language model assessment for reference structure generation, 3) qualitative comparison of relative vs.
+absolute alignment in LAGUNA’s features, 4) comparison with zero-shot MLLMs, 5) caption quality impact analysis, and 6) universal domain adaptation performance evaluation.
+Additional complementary ablation studies are provided in the supplementary material.
 Setting $\mathcal{L}_{S}$ $\mathcal{A}$ $\mathcal{A}_{t/s}$ CD Attn. $\mathcal{L}_{Reg}$ Avg. Acc. Rel. Imp. Abs. Imp.
@@
 Table 4, ablates the main methodological elements of LAGUNA on GeoImnet.
-We set up this ablation by removing LAGUNA’s core elements, namely the structure loss ($\mathcal{L}_{S}$ Eq.
+We set up this ablation by removing LAGUNA’s core elements, namely the structure loss ($\mathcal{L}_{S}$ - Eq.
 (5)), domain-agnostic reference anchors $\mathcal{A}$, learnable anchors $\mathcal{A}_{t/s}$, cross-domain attention (CD Attn.
 - Eq.
-(9), (10)), and the regularization loss ($\mathcal{L}_{Reg}$ Eq.
+(8), (9)), and the regularization loss ($\mathcal{L}_{Reg}$ - Eq.
 (7)).
@@
 first row).
-Absolute alignment In setting (1), the visual classifier is trained on source and target samples using labels and pseudo-labels predicted by the text supervisor, which leads to a baseline model akin to LaGTran [27].
-In setting (2), we add the reference anchors $\mathcal{A}$ and train the visual classifier as in setting (1), but we also add a loss to align visual representations with the related class anchors in absolute coordinates through cosine similarity.
-This yields an improvement of $+0.87$ with respect to (1), suggesting that imposing a reference structure beyond pseudo-labels is beneficial to performance and that absolute alignment is too restrictive.
-Relative alignment In setting (3), we extend (2) with a set of domain-independent learnable anchors ($\checkmark$*✓ ∗ in Table 4, indicating that the weights of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are shared) and add the structure-preserving loss $\mathcal{L}_{S}$ (Eq.
-6).
-The introduction of learnable anchors enables the relative alignment of the visual domain with the reference language domain, leading to an improvement of $+1.32$ with respect to (2) and $+2.10$ over (1), which highlights the benefits of the proposed relative alignment over an absolute one.
-Domain-Specific Anchors Adding domain-specific anchors in setting (4) allows source and target domains to focus on the individual characteristics of the respective domains, leading to a further improvement of $+1.77$ with respect to (3) and a robust $+3.87$ with respect to the baseline (1).
-Cross-Domain Attention Adding cross-domain attention in setting (5) bridges the gap between target and source representations, leading to a of $+0.44$ with respect to (4) and an overall improvement of $+4.31$ with respect to (1).
-Regularization Loss We achieve the final configuration in LAGUNA, where we add the regularization loss, which prevents collapse in learning domain-specific anchors and leads to a further improvement of $+1.16$ with respect to (5) and an overall gain of $+5.47$ with respect to (1).
-Together, these improvements show the effectiveness of LAGUNA, highlighting the benefits of learning domain-specific visual representations aligned to a reference language structure.
+Absolute alignment In setting (1), the visual classifier is trained on source and target samples using labels and pseudo-labels predicted by the text supervisor, which leads to a baseline model akin to LaGTran.
+In (2), we add $\mathcal{A}$ and train the visual classifier as in setting (1), but we also add a loss to align visual representations with the related class anchors in absolute coordinates through cosine similarity.
+Setting (2) improves (1) by +0.87+0.87, suggesting that imposing a reference structure beyond pseudo-labels is beneficial to performance and that absolute alignment is too restrictive.
+Relative alignment In (3), we extend (2) with domain-independent learnable anchors ($\checkmark*$ in Table 4, indicating that weights of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are shared) and add the $\mathcal{L}_{S}$ loss (Eq.
+5).
+The introduction of learnable anchors enables the relative alignment of the visual domain with $\mathcal{A}$, leading to gains of +1.32+1.32 w.r.t (2) and +2.10+2.10 over (1), highlighting the benefits of the proposed relative alignment.
+Domain-Specific Anchors Adding domain-specific anchors in setting (4) allows source and target domains to focus on the individual characteristics of the respective domains, leading to a further improvement of +1.77+1.77 w.r.t (3) and a robust +3.87+3.87 w.r.t the baseline (1).
+Cross-Domain Attention Adding cross-domain attention in setting (5) bridges the gap between target and source representations, leading to+0.44+0.44 w.r.t (4) and an overall improvement of +4.31+4.31 w.r.t (1).
+Regularization Loss We achieve the final configuration in LAGUNA, where we add the regularization loss, which prevents collapse in learning domain-specific anchors and leads to a further improvement of +1.16+1.16 w.r.t (5) and an overall gain of +5.47+5.47 w.r.t (1).
+These improvements show the efficacy of LAGUNA and the benefits of aligning visual representations to a reference structure.
+Figure 3: In a), similarity maps of $100$ randomly selected classes from GeoImnet (yellow for high similarity) and average accuracies.
+In b), t-SNE plots for 1000 randomly selected Source and Target samples from GeoImnet, with respective MMD scores in Relative (right) and absolute (left) spaces.
 Language models for reference structures.
 Fig.
-3 reports semantic similarity maps for $100$ randomly selected classes form GeoImnet when three different language models are used in stage 1 (SentenceTransformer [57], CLIP [56], and BERT [61]), together with the average accuracy obtained keeping stages 2-3 fixed.
-Notably, accuracy follows the ability of models to distinguish distinct classes, with low off-diagonal values indicating more discriminative models.
-While we achieve best results with SentenceTransformer, LAGUNA’s performance is robust to different reference structures, maintaining SOTA results also when the same language model is used across all stages (i.e., BERT).
-Figure 3: Similarity maps of $100$ randomly selected classes from GeoImnet (yellow for high similarity) and average accuracies.
-Ratio of target samples.
-LAGUNA demonstrates strong generalization capabilities even with limited target domain data (Table 5).
-We progressively increase the number of randomly selected pseudo-labeled target data used for training (10%-75%) and observe the impact on performance.
-Notably, LAGUNA achieves high accuracy even with only $10\%$ of the samples and attains SOTA results with just $20\%$ (compared with  [27]).
-Dataset 10% 20% 30% 50% 75% 100% [27]
-GeoImnet 61.40 63.43 64.21 66.42 66.71 68.68 63.92
-GeoPlaces 55.35 57.85 58.95 60.02 60.75 62.33 56.58
-Avg. 58.38 60.64 61.58 63.22 63.72 65.40 60.24
-Table 5: Average accuracy with different percentages of target data compared with previous SOTA LaGTran [27]. LAGUNA to obtains SOTA accuracy with $20\%$ (in green).
-Alignment in Absolute and Realtive Coordinates.
-Figure 4 shows t-SNE plots and related Maximum Mean Discrepancy (MMD) scores for $1000$ randomly selected samples from the GeoImnet validation set, where source is USA and target is Asia.
-Notably, features are separated in absolute coordinates, while they blend together (low MMD value) in relative coordinates thanks to the structure-preserving loss $\mathcal{L}_{S}$ and domain-specific learnable anchors.
+3 a) reports semantic similarity maps for $100$ randomly selected classes form GeoImnet when three different language models are used in stage 1 (SentenceTransformer, CLIP, and BERT), together with the average accuracy obtained keeping stages 2-3 fixed.
+Notably, accuracy follows the ability of models to distinguish classes, with low off-diagonal values indicating more discriminative models.
+While we achieve best results with SentenceTransformer, LAGUNA is robust to different reference structures, maintaining SOTA results.
+Absolute vs Relative Alignment.
+Figure 3 b) shows t-SNE plots and related Maximum Mean Discrepancy (MMD) scores for $1000$ randomly selected samples from the GeoImnet validation set, where the source is USA.
+Notably, features are separated in absolute coordinates, while they blend together (low MMD) in relative ones thanks to $\mathcal{L}_{S}$ and domain-specific learnable anchors.
 This highlights that semantic alignment can be achieved without forcing representations into a shared space, allowing each domain to preserve its unique characteristics while adhering to a common structure.
-Figure 4: t-SNE plots for 1000 randomly selected Source and Target samples from GeoImnet, with respective MMD [28] scores. Relative coordinates (right) are better aligned than absolute (left).
-Figure 5: Ablation with BLIP-1 and BLIP-2 generated captions.
+Comparison with ’Zero-Shot’ MLLMs.
+Despite recent advances in MLLMs with billions of parameters trained on web-scale data, our experiments reveal that LAGUNA significantly outperforms them in challenging domain adaptation scenarios.
+Testing on two benchmarks requiring fine-grained spatial and temporal discrimination (GeoNet and Ego2Exo, Tab 5&6), we compare against five SOTA 7-8B MLLMs.
+LAGUNA achieved +11.84% higher accuracy on GeoNet and +2.78% in mean per-class accuracy on Ego2Exo.
+More importantly, in the Ego2Exo benchmark, we surpass InternVL-8B, which is trained in egocentric data and holds the state-of-the-art in many egocentric benchmarks.
+As a final remark, LAGUNA is approximately 100× smaller than these MLLMs, demonstrating that tailored models are still better at challenging UDA tasks.
+Model Geolmnet GeoPlaces Avg.
 […]
-Computational Overhead.
-While LAGUNA involves three stages and different components, its training is quite efficient: Stage 1 needs no training, and embedding 600 class labels takes 1-2 minutes ($\sim$4GB GPU memory).
-On GeoImNet, with a V100, Stage 2 requires only 9 minutes per epoch, while Stage 3 adds $<$2 minutes per epoch, which is minor considering the baseline’s 65 minutes per epoch.
-At test time, LAGUNA only requires the visual encoder, the learned anchors, the light cross-attention layer and the final MLP classifier, bringing no noticeable overhead.
+LLaVA-Next 7B Li et al. (2024b) 22.88 24.21 40.41 42.33 32.46
+LLaVA-OneVision 7B Li et al. (2024a) 27.07 29.87 46.29 49.23 38.12
+Qwen2.5-VL 7B Bai et al. (2025) 52.98 56.34 50.85 54.11 53.56
+InternVL3 8B Zhu et al. (2025) 38.95 41.58 53.66 52.91 46.75
+PLM Cho et al. (2025) 19.93 23.96 35.58 42.61 30.52
+LAGUNA 67.39 69.97 61.15 63.51 65.40
+Improvement +14.41 +13.63 +7.49 +9.40 +11.84
+Table 5: Comparison on GeoImnet and GeoPlaces against zero-shot MLLMs.
+[…]
+LLaVA-Next 7B Li et al. (2024b) 4.28 4.95 4.62
+LLaVA-OneVision 7B Li et al. (2024a) 5.57 10.18 7.88
+Qwen2.5-VL 7B Bai et al. (2025) 11.46 20.39 15.93
+InternVL3 8B Zhu et al. (2025) 13.07 28.34 20.71
+PLM Cho et al. (2025) 22.54 37.16 29.85
+LAGUNA 13.52 33.45 23.49
+Improvement +0.45 +5.11 +2.78
+Table 6: Comparison on Ego2Exo against zero-shot MLLMs. PLMs training set includes EgoExo-4D.
+[…]
+LLaVA-OneVision 7B Li et al. (2024a) 82.20 74.10 78.15
+Qwen2.5-VL 7B Bai et al. (2025) 82.35 74.26 78.31
+LAGUNA (LLaVA captions) 82.51 76.20 79.35
+LAGUNA (Qwen2.5 captions) 82.77 76.19 79.48
+Table 7: Ablation on the quality of generated captions with different MLLMs on sketch and clipart domains from DomaiNet.
+Model H-Score
+Baseline Kalluri et al. (2024) 50.20
+UniDA You et al. (2019) 33.39
+DANCE Saito et al. (2020) 51.75
+OVANet Saito and Saenko (2021) 47.26
+LaGTran Kalluri et al. (2024) 61.19
+LAGUNA 64.41
+Table 8: Experiment on GeoUniDA for universal domain adaptation.
+Caption Quality.
+Table 1 presents LAGUNA results using BLIP-2-generated captions for consistency and fair comparison with previous works.
+In Table 7, we investigate how caption quality affects LAGUNA’s performance by generating captions from two SOTA MLLMs: LLaVA-OneVision 7B and Qwen2.5-VL 7B, and evaluating LAGUNA on DomainNet’s sketch and clipart domains, comparing its performance against the same MLLMs.
+Results show that LAGUNA benefits from higher-quality captioning models and offers an alternative that is 100x smaller while achieving better performance on specific domains.
+The supplementary material provides additional experiments demonstrating: 1) LAGUNA’s robustness to low-quality captions and 2) its requirement for only a few target captions to achieve SOTA performance, reducing the one-time computational cost of automatic annotation with MLLMs.
+These findings make LAGUNA well-suited for real-world deployment in constrained environments where billion-parameter models are impractical.
+Universal Domain Adaptation.
+Finally, we assess LAGUNA’s performance in open-world domain transfer scenarios using the GeoUniDA dataset Kalluri et al.
+(2023), which, in contrast with traditional domain adaptation benchmarks, evaluates universal domain adaptation across domains, including both unique and shared semantic classes.
+Following OVANet Saito and Saenko (2021), we employ the H-score metric that balances closed-set and open-set accuracies through their harmonic mean.
+Notably, LAGUNA exhibits strong performance in this challenging setting, validating the effectiveness of our method.
 5 Conclusions
-We introduced LAGUNA, a novel domain adaptation approach leveraging geometrical structures of semantically equivariant spaces to guide adaptation.
+We introduce LAGUNA, a novel domain adaptation approach leveraging geometrical structures of semantically equivariant spaces to guide adaptation.
 LAGUNA defines a reference representation space structure based on domain-agnostic class semantic similarities encoded by a language model, ensuring the organization of sample projections reflects this structure while preserving domain-specific characteristics.
@@
 This is achieved through pseudo-labeling and learnable domain-specific anchors, guided by a loss function that prioritizes mimicking geometrical associations over direct representation alignment.
-Extensive experiments demonstrate LAGUNA’s superior performance compared to existing methods.
-Supplementary Material
+Extensive experiments demonstrate LAGUNA’s superior performance compared to existing methods, highlighting the importance of structural alignment and language-guided learning in domain adaptation.
+Future work could explore extending LAGUNA to multi-modal adaptation scenarios.
+Appendix LAGUNA: LAnguage Guided UNsupervised Adaptation with structured spaces
+Figure 4: LAGUNA’s accuracy with different quantities of pseudo-labeled data (target samples with captions).
 Appendix A Introduction
-In this document, we report additional ablation studies regarding the choice of language models and the structural guidance of the target domain.
-Moreover, we provide a more detailed explanation of the motivation behind our regularization loss $\mathcal{L}_{Reg}$ in Eq.
+This supplementary material presents additional ablation studies complementing those in the main manuscript.
+We first examine the number of captions LAGUNA requires to achieve SOTA performance.
+Second, we extend caption quality experiments by testing LAGUNA’s robustness with lower-quality captions.
+Third, we ablate language model choices and target domain structural guidance.
+Fourth, we provide detailed motivation for our regularization loss $\mathcal{L}_{Reg}$ in Eq.
 (8).
-Finally, we attach a Python file, LAGUNA.py, demonstrating the implementation of LAGUNA in PyTorch.
-We will release the full code and configurations upon acceptance.
-Appendix B Ablation on Language Models
+Full code and implementation will be released upon acceptance.
+In addition to LAGUNA’s code, we will also provide scripts to reproduce all the MLLMs results reported in this work.
+Appendix B Ablation on the ratio of target samples.
+Our approach relies on image captions, whether generated or sourced elsewhere.
+A potential concern with LAGUNA is that automatic annotation of large datasets can be tricky to obtain in full or even time-consuming.
+While the costs of obtaining these captions are minimal compared to manual annotation, we investigate the quantity of captions LAGUNA requires for SOTA performance in Fig.
+4.
+We progressively increase randomly selected pseudo-labeled training data from 10% to 75% and observe the performance impact.
+Notably, LAGUNA achieves high accuracy with only 10% of samples and reaches SOTA results with just 20%.
+This efficiency highlights the benefits of our structure-driven approach, which organizes sample projections within domain-specific spaces according to a reference geometric structure.
+Appendix C Ablation on captioning quality.
+[…]
+Model S $\rightarrow$ C C $\rightarrow$ S
+BLIP-1 Li et al. (2022) 69.9 60.1
+BLIP-2 Li et al. (2023) 72.0 64.1
+LAGUNA (BLIP-1) 78.1 71.3
+LAGUNA (BLIP-2) 80.2 73.5
+Table 9: Ablation on caption quality with BLIP-1&2 captioning models. We consider sketch and clipart domains from DomainNet. The best result is in bold, and the second best is underlined.
+Appendix D Ablation on Language Models
 In addition to the language model used for defining the reference structure through $\mathcal{A}$ in Stage 1, LAGUNA employs language models in Stage 2 for encoding image/video captions trained for text classification and capturing semantic structure.
 This model is then used to generate pseudo-labels and serve as structure guidance for the target domain in Stage 3.
-We ablate the choice of language model in these Stages, comparing LAGUNA’s performance when employing SentenceTransformer [57], CLIP [56], or BERT [61] on GeoImnet dataset in two settings: (1) using SentenceTransformer-defined $\mathcal{A}$ (Table 6) as it can better model semantic relationships (recall Fig.
+We ablate the choice of language model in these Stages, comparing LAGUNA’s performance when employing SentenceTransformer Reimers and Gurevych (2019), CLIP Radford et al.
+(2021), or BERT Sanh et al.
+(2019) on GeoImnet dataset in two settings: (1) using SentenceTransformer-defined $\mathcal{A}$ (Table 10) as it can better model semantic relationships (recall Fig.
 3 from the main manuscript) and (2) defining $\mathcal{A}$ with the same language model as the one chosen for training (Table 11).
@@
 The best result is in bold, and the second best is underlined.
-Appendix C Ablation on structure guidance for target domain in Stage-3
+Appendix E Ablation on structure guidance for target domain in Stage-3
 During Stage-3 training of LAGUNA, in addition to classification training, we train our model also for structure-preserving as defined from $\mathcal{A}$ using the structural loss $\mathcal{L}_{S}$.
@@
 Note that for Ego2Exo the reported metric is mean per class accuracy.
-Appendix D Regularization Loss Extended
+Appendix F Regularization Loss Extended
 In Section 3.4.1 of the main manuscript, we explain how the structure loss $\mathcal{L}_{S}$ is supported by a regularization loss $\mathcal{L}_{Reg}$ to avoid feature collapsing.
@@
 where $\gamma=\mathcal{A}^{T}\mathcal{A}$.
+Appendix G Use of generative AI
+According to the ICLR 2026 policy regarding the transparency of generative AI use in submitted works, we declare that generative AI tools were employed solely for language enhancement and correction purposes in this manuscript.
+All content, including technical contributions, experimental design, analysis, and conclusions, represents original work by the authors.
+We take full responsibility for every word reported in this paper, as each statement has been carefully reviewed, verified, and deliberately selected to accurately represent our research findings and contributions.
+The use of AI assistance was limited exclusively to improving grammatical clarity and linguistic precision, without influencing the scientific content or integrity of our work.